### v1

# Inference Time Alignment of Diffusion Models with Evolutionary Algorithms

Inference Time Alignment of Diffusion Models with Evolutionary Algorithms
Purvish Jajal*
Purdue University
West Lafayette, IN 47907
pjajal@purdue.edu
&Nick John Eliopoulos*
Purdue University
West Lafayette, IN 47907
neliopou@purdue.edu
&Benjamin Shiue-Hal Chou
Purdue University
West Lafayette, IN 47907
chou150@purdue.edu
&George K. Thiruvathukal
Loyola University Chicago
Chicago, IL 60660
gkt@cs.luc.edu
&James C. Davis
Purdue University
West Lafayette, IN 47907
davisjam@purdue.edu
&Yung-Hsiang Lu
Purdue University
West Lafayette, IN 47907
yunglu@purdue.edu
Evolutionary Algorithms for Inference-Time Alignment of Diffusion Models
Purvish Jajal*
Purdue University
West Lafayette, IN 47907
pjajal@purdue.edu
&Nick John Eliopoulos*
Purdue University
West Lafayette, IN 47907
neliopou@purdue.edu
&Benjamin Shiue-Hal Chou
Purdue University
West Lafayette, IN 47907
chou150@purdue.edu
&George K. Thiruvathukal
Loyola University Chicago
Chicago, IL 60660
gkt@cs.luc.edu
&James C. Davis
Purdue University
West Lafayette, IN 47907
davisjam@purdue.edu
&Yung-Hsiang Lu
Purdue University
West Lafayette, IN 47907
yunglu@purdue.edu
Evolutionary Algorithms for Black-Box Inference-Time Alignment of Diffusion Models
Purvish Jajal*
Purdue University
West Lafayette, IN 47907
pjajal@purdue.edu
&Nick John Eliopoulos*
Purdue University
West Lafayette, IN 47907
neliopou@purdue.edu
&Benjamin Shiue-Hal Chou
Purdue University
West Lafayette, IN 47907
chou150@purdue.edu
&George K. Thiruvathukal
Loyola University Chicago
Chicago, IL 60660
gkt@cs.luc.edu
&James C. Davis
Purdue University
West Lafayette, IN 47907
davisjam@purdue.edu
&Yung-Hsiang Lu
Purdue University
West Lafayette, IN 47907
yunglu@purdue.edu
Inference-Time Alignment of Diffusion Models with Evolutionary Algorithms
Purvish Jajal*
Purdue University
West Lafayette, IN 47907
pjajal@purdue.edu
&Nick John Eliopoulos*
Purdue University
West Lafayette, IN 47907
neliopou@purdue.edu
&Benjamin Shiue-Hal Chou
Purdue University
West Lafayette, IN 47907
chou150@purdue.edu
&George K. Thiruvathukal
Loyola University Chicago
Chicago, IL 60660
gkt@cs.luc.edu
&James C. Davis
Purdue University
West Lafayette, IN 47907
davisjam@purdue.edu
&Yung-Hsiang Lu
Purdue University
West Lafayette, IN 47907
yunglu@purdue.edu
Abstract
Diffusion models are state-of-the-art generative models in various domains, yet their samples often fail to satisfy downstream objectives such as safety constraints or domain-specific validity. Existing techniques for alignment require gradients, internal model access, or large computational budgets. We introduce an inference-time alignment framework based on evolutionary algorithms. We treat diffusion models as black-boxes and search their latent space to maximize alignment objectives. Our method enables efficient inference-time alignment for both differentiable and non-differentiable alignment objectives across a range of diffusion models. On the DrawBench and Open Image Preferences benchmark, our EA methods outperform state-of-the-art gradient-based and gradient-free inference-time methods. In terms of memory consumption, we require $55\%$ to $76\%$ lower GPU memory than gradient-based methods. In terms of running-time, we are $72\%$ to $80\%$ faster than gradient-based methods. We achieve higher alignment scores over 50 optimization steps on Open Image Preferences than gradient-based and gradient-free methods.
footnotetext: * Purvish Jajal and Nick John Eliopoulos contributed equally to this work.
1 Introduction
Figure 1: Samples generated by our method on Stable Diffusion-3 across three different prompts under various alignment objectives (brightness, saturation, file size). Each row shows the progression over optimization steps $i$, with the corresponding metric values displayed in the top-right corner. The noise in each generation is optimized via the CoSyNE algorithm [13] over 14 optimization steps. Arrows ($\uparrow$/$\downarrow$) indicate whether the metric is being maximized or minimized.
Diffusion models [40, 34, 9] have emerged as state-of-the-art generative models for synthesizing high-quality images, video, and audio [3, 16, 18, 25]. However, the outputs of diffusion models often fail to meet downstream objectives, such as adhering to user preferences [53, 51, 37], safety constraints [28], or domain-specific validity [14, 49]. This problem is referred to as alignment—ensuring that generated samples satisfy external objectives beyond the original maximum-likelihood objective [24, 48, 2, 21, 5, 31, 4]. Alignment is an important concern for many practical applications such as image generation [53, 51, 37] and drug discovery [14, 49].
Methods for diffusion model alignment can be categorized into two types: fine-tuning-based alignment and inference-time alignment. Fine-tuning-based alignment typically relies on supervised fine-tuning or reinforcement learning to adjust model parameters [24]. In contrast, inference-time alignment modifies the sampling process during generation to maximize a reward function (i.e., alignment objective) [45]. These methods aim to be computationally efficient and support arbitrary rewards while maintaining sample quality. However, existing approaches have several drawbacks: they are often not black-box, requiring access to gradients or internal model states; they cannot efficiently generate many aligned samples; and they are often not sample-efficient. As a result, they are difficult to integrate with a variety of models, sampling strategies, and reward functions; generating many aligned samples becomes cost-prohibitive; and obtaining aligned samples may take too long.
In this work, we demonstrate that evolutionary algorithms (EAs)—specifically genetic algorithms (GA) and evolutionary strategies (ES)—can be used to align diffusion models. As a class of algorithms, EAs have desirable properties: they are black-box, gradient-free optimizers, that can be highly parallelized [39, 36]. By interpreting the diffusion models as black-boxes and searching over their latent space by using EAs we can optimize the alignment objective. Our approach enables efficient alignment that supports both differentiable and non-differentiable rewards, and is compatible with various diffusion models. Fig. 1 depicts our method applied to various alignment objectives.
To quantify alignment quality, GPU-memory footprint, and runtime efficiency, we evaluate our inference-time alignment method on two prompt collections [35, 6], four reward functions [53, 51, 32, 4], and three diffusion models [34, 9, 3]. On both prompt collections, our method outperforms existing gradient-free and gradient-based inference-time alignment approaches. In terms of computational costs, we exhibit $55\%$ to $76\%$ reduction in GPU memory than gradient-based methods, while only consuming up to 1.5×1.5$\times$1.5 × more memory than gradient-free methods. When producing the same number of aligned samples, evolutionary methods achieve 3.5×3.5$\times$3.5 × to $\times5$ lower runtime per optimization step compared to gradient-based methods, and only $\times1$ to 1.6×1.6$\times$1.6 × higher running time per step than gradient-free methods. Moreover, evolutionary methods reach higher alignment scores than gradient-free and gradient-based methods in 10 optimization steps. In sum, we contribute:
  • We propose a novel inference-time alignment approach using evolutionary algorithms (EAs). Specifically, we introduce two black-box methods for aligning the outputs of diffusion models: (1) optimizing noise, and (2) optimizing transformations applied to the noise. We elaborate upon the requisite choices for utilizing EAs for inference-time alignment, such as search algorithms, initialization, operators, and computational considerations.
  • We demonstrate the effectiveness of two widely used classes of evolutionary algorithms on the diffusion model alignment task: genetic algorithms and evolutionary strategies. Our evaluation spans a diverse set of reward functions and diffusion models.
2 Related Work
In this section, we review prior work on diffusion model alignment (§ 2.1) and evolutionary algorithms (§ 2.2). An extended treatment of related works can be found in § B.1.
2.1 Diffusion Model Alignment
Diffusion models use a reverse diffusion process to convert some latent noise distribution into a data distribution, such as images [40, 16]. The reverse diffusion process iteratively denoises the initial latent noise $z_{T}$ over some number of steps $T$ to yield a sample, $z_{0}$. Though diffusion models are capable of modeling complex data distributions, they often fail to produce samples that meet some downstream objective. Diffusion model alignment methods adjust diffusion models such that the resulting samples better meet an objective beyond the model’s original maximum likelihood criterion. These objectives commonly focus on producing images that reflect human aesthetic preferences [17, 51, 52, 53], or other metrics such as compressibility [2].
Alignment methods can generally be grouped into two main categories: fine-tuning–based methods and inference-time methods. Fine-tuning-based alignment methods involve adjusting the diffusion model’s parameters so that generated samples better match alignment objectives [21, 5, 31, 4, 48, 2]. Fine-tuning based methods, although powerful, require retraining and thus all the associated costs.
We follow the alternative approach, inference-time methods. Rather than altering model parameters, these techniques adjust sampling or conditioning to ensure samples meet alignment objectives. Formally, they seek the control variable $\psi$ that maximizes the expected reward $R(x)$ of samples $x$ from a pretrained diffusion model $p_{\theta}$, as shown in Eq. 1.
$\psi^{\star}=\underset{\psi\in\Psi}{\arg\max}\;\mathbb{E}_{x\sim p_{\theta}(x\mid\psi)}\bigl{[}R(x)\bigr{]}.$ (1)
Examples of control variables, $\psi$, include optimizing conditional input prompts [15, 12, 27, 10], manipulating cross-attention layers [11], and tuning latent noise vectors (noise optimization) to guide the diffusion trajectory [47, 42, 23, 45, 55, 26]. In this work, we focus exclusively on noise optimization ($\psi=z$), which is the most general method. Broadly, noise optimization methods fall into gradient-based optimization, or gradient-free optimization. We discuss each approach in turn.
Gradient-based methods refine the noise iteratively by leveraging the gradient with respect to the reward. Direct Optimization of Diffusion Latents (DOODL) [47] and Direct Noise Optimization (DNO) [42] are exemplary of these approaches. Both need to contend with the computational costs of backpropagation and maintaining sample quality by keeping the optimized noise on the Gaussian shell—noise sampled from high-dimensional Gaussians concentrate on the surface of a spherical shell [46]. The drawbacks of gradient-based methods are: they are not black-box; they are computationally expensive, when aligning multiple samples; and require long optimization budgets.
Gradient-free methods, on the other hand, explore the space of noise vectors or trajectories using only search or sampling methods. Uehara et al. [45] provide a comprehensive overview of various inference-time algorithms, largely covering sampling-based methods. Ma et al. [26] employ three different sample search strategies — random search, zero-order search (similar to hill climbing), and search-over-paths. Li et al. propose DSearch [23], a dynamic beam search algorithm. These techniques vary in their trade-offs. DSearch and the techniques outlined by Uehara et al. are not black-box and may have large computational costs. Although random and zero-order search are also black-box and computationally efficient, evolutionary methods outperform them (§ 4.2).
2.2 Evolutionary Algorithms
Evolutionary algorithms (EAs) are a class of biologically inspired methods that can be used to solve a variety of black-box optimization problems [8, 39]. They iteratively select, recombine, and mutate populations of solutions to improve their fitness (quality of the solution). Among the various EAs, genetic algorithms (GAs) [8, 39] and natural evolutionary strategies (ES) [50] are most pertinent to this work. Genetic algorithms and natural evolutionary strategies have been widely applied in deep learning and deep reinforcement learning. In deep learning, genetic algorithms have facilitated neural architecture search and hyperparameter tuning [33, 29, 54]. In deep reinforcement learning, both genetic algorithms and natural evolutionary strategies have been shown to be competitive with back-propagation based methods such as policy gradients [41, 36]. GA and ES are also used in image generation [22, 38, 43, 10] and artificial life simulations [20]. However, to our knowledge, we are the first work to apply EAs for black-box diffusion model alignment.
3 Inference Time Alignment via Evolutionary Algorithms
We cast inference-time alignment as a black-box search problem and apply an evolutionary search framework. For a pretrained diffusion model $p_{\theta}$ and reward function $R(x)$, we aim to find parameters $\phi$ of the search distribution $q_{\phi}(\psi)$ that maximizes the expected reward in Eq. 2.
$\phi^{\star}=\arg\max_{\phi}\;\underbrace{\mathbb{E}_{\psi\sim q_{\phi}(\psi)}}_{\begin{subarray}{c}\text{search}\\ \text{distribution}\end{subarray}}\Bigl{[}\mathbb{E}_{x\sim p_{\theta}(x\mid\psi)}\bigl{[}R(x)\bigr{]}\Bigr{]}\,.$ (2)
For generality, we write $x\sim p_{\theta}(x\mid\psi)$ due to many diffusion samplers being stochastic. In practice, we treat this as a deterministic map from control variable to sample, i.e., $f_{\theta}:\psi\mapsto x$. See § B.2 for further details.
In § 3.1, we review design goals and requirements for our method. In § 3.2 we define the solution spaces ($\psi$), over both noise and noise-transformations. In § 3.3 we describe how genetic algorithms and natural evolutionary strategies define the search distribution $q_{\phi}(\psi)$ and how they optimize $\phi$ in Eq. 2. Finally, § 3.4 presents our implementation and computational considerations.
3.1 Design Criteria
Prior works [45, 42] have outlined a set of Criteria that characterize effective inference-time alignment approaches: (C1) maintaining sample quality; (C2) computationally efficient; and (C3) supporting non-differentiable reward functions. In this work, we advocate a fourth criterion: (C4) alignment methods should strive to be black-box. Black-box inference-time alignment can be easily integrated with a variety of models, sampling strategies, and reward functions. For example, a black-box alignment method could be applied to a quantized diffusion model for fast inference, and could simultaneously support arbitrary reward functions.
Figure 2: Illustration of our evolutionary search approach to noise vector optimization. Here, we depict the connection between terms of Eq. 2 and components of our method via color-coding.
3.2 Defining the Solution Space
Before applying evolutionary algorithms, we must define the solution space, i.e., what we will search over, $\psi$. For aligning diffusion models, we explore two options: (1) directly searching over noise vectors, and (2) searching over transformations of an initial noise vector. In both cases, our overarching goal is to identify a solution that maximizes Eq. 2.
Searching over Noise. The simplest solution space we can define is over the initial noise, i.e., $\psi=z_{T}^{\prime}$ in Eq. 2. Concretely, we define the search variable $\psi$ to be the initial noise vector, $z_{T}^{\prime}\sim q_{\phi}(z_{T}^{\prime})$, and optimize $\phi$ to maximize the expected reward. This direct search is straightforward, but hinges on proper initialization and careful tuning of mutation; without these, evolved solutions can drift into low-density regions of the latent space, leading to poor sample quality.
Searching for Noise Transformations. Instead of searching over noise vectors directly, we can define an affine transformation of an initial noise vector ($z_{T}$) and search over its parameters. We define the transformed noise as $z_{T}^{\prime}=f(z_{T})=Az_{T}+b$, where $A\in\mathbb{R}^{d\times d}$, $z_{T},b\in\mathbb{R}^{d}$. Thus, in Eq. 2 we set $\psi=[A,b]$ with $[A,b]\sim q_{\phi}([A,b])$ and optimize $\phi$ to maximize the expected reward in Eq. 2 under $x\sim p_{\theta}(x\mid Az_{T}+b)$.
[…]
Choosing between noise and noise transformation search. The choice between noise and noise-transformation search hinges on the type of regularization each imposes and their relative computational costs. Noise search, while being the most direct and straightforward way to define our search space, does not confine solutions to the high-density shell of the Gaussian. Consequently, if the search algorithms (§ 3.3) fail to ensure shell-confinement, sample quality degrades. In contrast, searching over noise transformations will guarantee (given $b=0$, see § 3.2) residence within the high-density shell regardless of the choice of evolutionary algorithm. However, noise-transformation search incurs extra cost for factorization and linear transforms; we describe mitigation strategies in § 3.4.
3.3 Searching the Solution Space
Given our solution space ($\psi$) we now turn to how we define the search distribution and find its parameters, i.e., $q_{\phi}(\psi)$. Our requirements (C1C4) call for black-box, gradient-free optimizers that preserve sample quality and support arbitrary rewards. Among the large family of optimization approaches [7, 19, 8], evolutionary algorithms (EAs) satisfy these criteria and have a history of efficacy in machine learning contexts (§ 2.2).
We consider two families of EA: genetic algorithms (GAs) and natural evolutionary strategies (ES). For each, we first review their core mechanics, then describe how these mechanics map to solving the diffusion alignment task. Their specific optimization procedures are outlined in § B.3.
Genetic Algorithms (GAs). Genetic algorithms maintain a population of candidate solutions, $\psi_{i}$, and iteratively improve them via selection, crossover, and mutation [8]. In Eq. 2, the search distribution is an empirical distribution over the population of solutions.
In the context of alignment, GAs satisfy the criteria in § 3.1 and offer practical advantages that improve alignment performance. First, the optimization procedure—selection, crossover, mutation—under correct parameterization can ensure the noise search is regularized to the high-density shell, provided the initial population resides within. In particular, uniform crossover (independent coordinate swaps) or permutation-based mutations do so—see § B.4 for details. Second, maintaining a population enables broad, parallel exploration early in optimization, while increasing selection pressure later drives convergence towards the promising regions (§ D.4). This capacity to explore multiple basins simultaneously makes GAs well-suited to multi-modal alignment objectives, e.g., human preferences. However, our experiments show that diversity collapse can cause premature convergence, hence marginal improvements over longer optimization timelines (see § 4.3 and § 4.4).
Natural Evolutionary Strategies (ES). Natural evolutionary strategies perform black-box optimization by adapting the parameterized search distribution $q_{\phi}$ over solutions [50]. Unlike GAs, $q_{\phi}$ is a parameterized search distribution (e.g., multivariate normal).
In the context of alignment, ES satisfy the criteria in § 3.1 and offer a distinct set of advantages over GAs. They maintain a search distribution, so their memory footprint is lower: only distribution parameters (e.g., mean, covariance, step size) need be stored. This parameterized distribution also lets us draw an arbitrary number of candidate solutions, making ES ideal when many aligned samples are required. Moreover, although termed “gradient-free”, evolutionary strategies update their search distribution with a gradient approximation [36, 41]. As a result, they move through the reward landscape with a coarse sense of direction lacked by GAs. This directional information translates into alignment improvements over long optimization horizons (see § 4.3).
Choosing Between GA and ES for Alignment. Although GA and ES methods satisfy the criteria outlined in § 3.1, they exhibit markedly different properties. First, the search distribution for GA is an empirical distribution, whereas ES use multivariate normal distributions. Consequently, GA exhibit higher memory usage than ES because the entire population must reside in memory (see § 4.5). This has consequences for sampling as well: additional samples outside the population cannot be acquired for GA, whereas for ES, it merely requires sampling the parameterized distribution. Conversely, GA handle multi-modal rewards better than ES due to the initialization requirements of each EA (§ 3.4). However, we note that GAs often suffer from premature convergence whereas ESs do not (§ 4.4).
3.4 Implementation and Computational Considerations
This section discusses practical considerations for implementing inference-time alignment using evolutionary algorithms, focusing on initialization strategies and parallel evaluation.
Initialization and Sample Quality. Initialization is essential for effective search. For GAs, initializing the population from the standard Gaussian distribution $\mathcal{N}(0,I)$ ensures that the initial samples reside within the high-density shell of the latent space. However, this initialization strategy does not translate to natural evolutionary strategies (ES), as subsequent updates to the search distribution can degrade sample quality. To mitigate this, we instead initialize $\mu$ of the search distribution to a fixed latent vector $z_{0}\sim\mathcal{N}(0,I)$ and use a small initial standard deviation. We elaborate further on consequences of ES initialization in Appendix C.
Parallel Evaluation and Population Scaling. Both GA and ES approaches evaluate a population of $N$ candidate solutions at each iteration. To improve runtime efficiency, we partition each population into batches of size $B$ and evaluate the batched solutions in parallel. Batching takes advantage of modern GPU parallelism and allows us to scale evaluations across larger populations without incurring excessive runtime costs. However, this comes at the cost of higher memory usage during model inference. In practice, the ability to choose the batch size for evaluation allows for a trade-off between memory usage and runtime. We explore this further in § 4.5.
4 Evaluation
We evaluate our noise optimization design decisions, and compare with other inference-time alignment work. § 4.1 describes our experimental setup. § 4.2 compares the ability of our method to maximize rewards with other work, notably Direct Noise Optimization [26] on the DrawBench and Open Image Preferences datasets. § 4.3 investigates the relationship between optimization steps and alignment. § 4.4 compares population statistics between ES and GA methods. § 4.5 reports the memory consumption and runtime of our method across population size and batch size.
4.1 Experimental Setup
We evaluate our alignment methods on a variety of diffusion models, benchmarks, and evolutionary algorithms. All experiments are conducted on NVIDIA A100 80GB GPUs.
Evolutionary Algorithms We use a variety of state-of-the-art evolutionary algorithms: CoSyne, PGPE, and SNES, as provided in EvoTorch [44].
Diffusion Models We implement EA-based search on three state-of-the-art diffusion models: Stable Diffusion-1.5 [34], Stable Diffusion 3 [9], and the latent consistency variant (LCM) of PixArt-$\alpha$ [3]. Our choices reflect the diversity of diffusion model architectures and training methods.
Datasets We follow [26], and conduct quantitative evaluations on DrawBench [35]. DrawBench consists of 200 prompts from 11 categories that test the semantic properties of model, such as their ability to handle composition, cardinality, and rare prompts. We conduct our ablation studies and qualitative evaluations on a subset (60 prompts, 5 per category) of the Open Image Preferences dataset [1]. Open Image Preferences consist of detailed and diverse prompts along with human preferences, which can be used to fine-tune diffusion models [6]. For our purposes, both datasets provide complex prompts to evaluate the quality of inference-time methods.
Reward Functions In accordance with other diffusion works [42, 23], we use ImageReward [53], CLIP [32], HPSv2 [51], and JPEG compressibility as reward functions and evaluation metrics.
4.2 Cross-Dataset Evaluation
DrawBench. Tab. 1 depicts the efficacy of EAs at alignment. After only 15 optimization steps, EAs outperform both baseline black-box methods (Random, Zero-Order) and the gradient-based DNO on all four reward functions, for both the best sample and the median reward in the population. When searching over noise, CoSyNE delivers the strongest overall performance, achieving the best alignment performance on ImageReward, CLIP, HPSv2 scores and yields the largest JPEG-size reduction. The same pattern holds when we search over noise transformations, with CoSyNE again leading three of the four metrics; the sole exception is the JPEG-compression objective, where PGPE attains the smallest file sizes. Manual inspection, however, reveals that many of PGPE’s images omit key prompt content, suggesting reward hacking (see § D.5.1 for examples). Crucially, both EAs and the Zero-Order baseline improve median as well as best rewards, whereas Random search does not — indicating that directed search can steer us toward parts of the latent space where aligned samples reside. We elaborate on the consequences of searchable latent spaces in § 5.
Table 1: DrawBench Evaluation on StableDiffusion-1.5 We show the best-sample reward and median reward of the best population achieved by each algorithm on DrawBench under four reward functions. Algorithms (Random, Zero-Order, DNO) are baselines, while (CoSyNE, SNES, PGPE) are evolutionary methods. We conduct search using each algorithm for 15 steps with a population of 16 solutions, except for DNO which has population of 1. NB: $\uparrow$/$\downarrow$ notation indicates a goal of reward maximization or minimization, respectively. Entries marked with a † are statistically insignificant ($p$ > 0.05), otherwise they are statistically significant.
Algorithm ($\uparrow$) ImageReward ($\uparrow$) CLIP ($\uparrow$) HPSv2 ($\downarrow$) JPEG Size (kB)
Best Median Best Median Best Median Best Median
Random $1.41$ $0.38$ $37.1$ $32.9$ $0.301$ $0.282$ $66.9$ $105.4$
Zero-Order $1.46$ $0.97$ $37.6$ $34.5$ $0.304$ $0.290$ $53.3$ $62.8$
DNO $0.71$ $26.2$ $0.286$ $94.6$
Ours: Noise
CoSyNE $\mathbf{1.61}$ $\mathbf{1.38}$ $\mathbf{38.8}$ $\mathbf{36.5}$ $\mathbf{0.310}$ $\mathbf{0.300}$ 39.5 44.8
SNES $1.39$ $0.78$ $38.1$ $35.6$ $0.300$ $0.286$ $71.3$ $78.0$
PGPE $1.26$ $0.92$ $37.1$ $34.4$ $0.299$ $0.290$ $88.3$ $97.0$
Ours: Noise Transformation (QR)
CoSyNE $\mathbf{1.53}$ $\mathbf{1.23}$ $\mathbf{38.4}$ $\mathbf{36.5}$ $\mathbf{0.307}$ $\mathbf{0.299}$ $38.7$ $42.3$
SNES $1.34$ $0.94$ $37.4$ $34.7$ $0.300$ $0.290$ $57.8$ $66.1$
PGPE $1.28$ $0.88$ $36.9$ $34.8$ $0.300$ $0.290$ $\mathbf{18.3}$ $\mathbf{20.2}$
Table 2: Open Images Preferences Cross-Model Evaluation The best-sample reward and the median reward of the best population achieved by each algorithm on Open Image Preferences across three models: StableDiffusion 1.5, StableDiffusion 3, and PixArt-$\alpha$. Algorithms are configured in the same manner as in  Tab. 1. Higher rewards are better.
Algorithm SD1.5 SD3 PixArt-$\alpha$ (LCM)
CLIP ImgReward CLIP ImgReward CLIP ImgReward
Best Med. Best Med. Best Med. Best Med. Best Med. Best Med.
Random $39.1$ $34.9$ $1.49$ $0.59$ $39.0$ $36.0$ $1.72$ $1.38$ $38.8$ $35.9$ $1.67$ $1.25$
Zero–Order $39.5$ $36.4$ $1.54$ $0.98$ $39.4$ $37.4$ $1.75$ $1.60$ $38.9$ $36.1$ $1.67$ $1.29$
Ours: Noise
CoSyNE $\mathbf{40.7}$ $\mathbf{38.3}$ $\mathbf{1.69}$ $\mathbf{1.47}$ $\mathbf{40.6}$ $\mathbf{39.0}$ $\mathbf{1.82}$ $\mathbf{1.75}$ $\mathbf{39.2}$ $\mathbf{36.8}$ $\mathbf{1.69}$ $\mathbf{1.45}$
SNES $40.1$ $37.6$ $1.57$ $1.25$ $39.9$ $38.1$ $1.77$ $1.68$ $38.9$ $35.8$ $1.68$ $1.21$
PGPE $39.0$ $36.1$ $1.37$ $0.88$ $38.4$ $36.4$ $1.67$ $1.50$ $38.7$ $35.8$ $1.65$ $1.23$
Open Image Preferences. Tab. 2 analyzes three diffusion models and confirms that evolutionary algorithms largely remain competitive across models. For both SD1.5 and SD3, the EA family consistently surpasses the black-box baselines: CoSyNE achieves the highest best and median rewards on every metric, e.g., SD3-CLIP rises from 39.4/37.4 with Zero-Order to 40.6/39.0, while SD3-ImageReward improves from 1.75/1.60 to 1.82/1.75. SNES and PGPE follow the same trend, but with smaller margins. Generalization is weaker on PixArt-$\alpha$, where improvements vanish, suggesting that black-box latent search may not transfer to architectures trained with alternate objectives such as LCM — we elaborate in § 5. Overall, we show evolutionary search can align multiple diffusion models; however some latent spaces may be more difficult to search than others.
4.3 Optimization Steps and Alignment
We evaluate how the number of inference-time optimization steps affects alignment. Fig. 3 depicts the best-sample and median reward per optimization step on Open Image Preferences benchmark, using ImageReward alignment objective.
All methods improve their best-sample and median rewards as the number of steps increases, but they differ in step-efficiency. The population-based EA Cosyne achieves the highest step-efficiency, yielding the highest best-sample and median reward at every step count. Whereas, the distribution-based EAs, SNES and PGPE, require more steps to surpass random and zero-order search. SNES needs more 15 than steps to outperform zero-order in best-sample performance. Meanwhile, PGPE needs more than 10 steps to exceed random search in best-sample performance. When considering median reward, all EAs exceed baselines given enough optimization steps.
(a) Best-sample reward per step on ImageReward.
(b) Median reward per step on ImageReward.
Figure 3: Reward per step measured on Open Image Preferences. Evolutionary methods (Cosyne, SNES, PGPE) outperform Zero-Order and Random baselines. Each algorithm optimizes the reward for 50 steps with a population of 16, except for DNO which has a population of 1.
4.4 Population Statistics
Fig. 4 plots the population’s reward standard deviation over 50 optimization steps for CoSyNE (a GA) and SNES (an ES), confirming the dynamics predicted in § 3.3. CoSyNE begins with high variance, reflecting broad exploration, but diversity collapses under selection pressure (§ D.4 gives ablations on selection pressures).
Figure 4: Reward standard deviation per step, measured on Open Image Preferences. CoSyne (GA) has high diversity early, but this rapidly diminishes. SNES (ES) has a more uniform pattern, as its search distribution adapts per step (§ 3.3). NB: Reward standard deviation proxies for solution diversity.
In contrast, SNES sustains a moderate level of variance throughout optimization, preserving exploratory capacity even at later steps. This sustained diversity underpins the long-horizon gains of ES (Fig. 3(a)). Taken together, these results suggest that GAs are ideal for rapid, short-horizon alignment, while ES are preferable when more time is available.
4.5 Computational Costs
We compare the memory usage and runtime per step for evolutionary methods in § 3 with Random, Zero-Order Search, and DNO. We measure the runtime per optimization step and the peak GPU memory usage with respect to batch size ($B$) and population size ($P$), as stated in § 3.4.
Fig. 5(a) shows that as batch size $B$ increases, evolutionary methods consume significantly less GPU memory than DNO. The gradient-based DNO consumes $\times2$ to 4.1×4.1$\times$4.1 × more GPU memory than evolutionary methods. Moreover, DNO runs out of memory (OOM) at batch sizes $\geq 16$, whereas evolutionary methods do not. Compared to other gradient-free approaches (Random and Zero-Order search), evolutionary methods consume only 1×1$\times$1 × to 1.5×1.5$\times$1.5 × more GPU memory. However, in terms of alignment performance (Tab. 1), evolutionary methods outperform both Random and Zero-Order search with small additional memory costs.
Fig. 5(b) shows the running time per step for each method as we vary the batch size $B$. All evolutionary algorithms (CoSyne, PGPE, SNES) exhibit a decrease in runtime per-step as $B$ grows: at small $B$, the population of $P$ candidates must be evaluated sequentially, thus runtime is high. As $B$ increases, evaluations are batched and parallelized, reducing the time per step at the expense of additional memory usage as shown in Fig. 5(a). In contrast, DNO’s per-step runtime increases with B and eventually exceeds all methods at large batch sizes.
(a) Memory usage vs. batch size ($P$= $B$).
(b) Running time per step vs. batch size. $(P=16)$
Figure 5: Computational characteristics of inference-time alignment methods on Stable Diffusion-1.5. In terms of memory usage evolutionary methods (CoSyne, PGPE, SNES) scale with batch size, whereas DNO does not and exhausts memory for batch sizes $\geq 16$. All methods aside from DNO exhibit near-constant or decreasing per-step latency as batch size increases. For DNO: $(P=B)$
5 Discussion and Limitations
Searching Latent Spaces. By virtue of our results, we show that optimizing noise with EAs is often sufficient to perform alignment (see Tab. 1), though its effectiveness varies by model (Tab. 2). This raises two questions: (1) Are some latent spaces easier to search than others? (2) How can one design, train, or modify diffusion models to make their latent spaces easier to search? Our reward improvements on PixArt-$\alpha$ were diminished compared to other diffusion models (Tab. 2), suggesting some property of the latent space and/or diffusion model that affects our search efficacy.
Limitations. Our approach is ill-suited for long optimization horizons. With enough time, white-box methods that exploit gradient information will achieve better alignment (see § D.2). Black-box alignment fares better on short time-horizon applications, e.g., resource-constrained tasks such as mobile image generation. It is less suitable in contexts with high resource budgets, e.g., drug discovery. Even over short optimization horizons, our EA techniques shares the limit of other gradient-free and gradient-based methods (§ 2.1): some engineering time is incurred in tuning, and regularization is required to maintain sample quality (§ 3). We did not evaluate our method across modalities other than images, but we conjecture that our method could perform well across modalities.
6 Conclusion
We illustrate how inference-time alignment for diffusion models can be performed with evolutionary algorithms (EA). Specifically, we search for the initial noise vector used in the reverse denoising process in order to alter the sampled image and maximize some reward. Our results illustrate that EAs can achieve higher rewards than existing inference-time alignment methods across a set of various diffusion models and datasets. In terms of compute resource requirements, our EA-based method uses 55%-76% lower GPU memory and runs 72%-80% faster than gradient-based alignment methods. Our findings highlight the effectiveness of EAs for inference-time alignment via noise optimization.
Appendix A Overview of Appendices and Supplemental Material
  • Appendix A: This summary.
  • Appendix B: Extended commentary on related work § 2.1 and § 2.2, with insights and details on our implementation of evolutionary algorithms for alignment.
  • Appendix C: Summary of insights we encountered when deciding how to initialize genetic algorithms and evolutionary strategies, related to § 3.3.
  • Appendix D: Extended quantitative and qualitative results across various models, reward functions, datasets, and algorithms used in our work, complementary to § 4.2§ 4.4.
Appendix B Extended Commentary on Related Work and Implementations
Here, we provide additional detail for concepts in our work that was not able to be fit into the main text.
B.1 Extended Diffusion Model Alignment Related Work
Diffusion models use a reverse diffusion process to convert some latent noise distribution into a data distribution, such as images [40, 16]. The reverse diffusion process iteratively denoises the initial latent noise $z_{T}$ over some number of steps ($t=T\xrightarrow{}t=0$) to yield a sample, $z_{0}$.
Though diffusion models are capable of modeling complex data distributions, they often fail to produce samples that meet some downstream objective. Diffusion model alignment methods adjust diffusion models such that the resulting samples better meet an objective beyond the model’s original maximum likelihood criterion. These objectives commonly focus on producing images that reflect human aesthetic preferences [17, 51, 52, 53], or other metrics such as compressibility [2].
Table 3: Comparison of inference-time diffusion alignment methods.
Method Arbitrary Rewards Gradient Free Black-box
DOODL [47]
DNO [42]
DSearch [23]
Search-Paths [26]
Zero-Order [26]
Ours
Alignment methods can generally be grouped into two main categories: fine-tuning–based methods and inference-time methods. Fine-tuning-based alignment methods involve adjusting the diffusion model’s parameters so that generated samples better match alignment objectives. Generally, these method rely on curated datasets [37] or reward models [53, 52, 51] to fine-tune diffusion models [21, 5, 31, 4], and can involve supervised fine-tuning, or reinforcement learning based methods [48, 2]. Fine-tuning based methods, although powerful, require retraining and thus all the associated costs — our work does not require training.
Our work follows the alternative approach, inference-time methods. Rather than altering model parameters, these techniques adjust sampling or conditioning to ensure samples meet alignment objectives. Formally, they seek the control variable $\psi$ that maximizes the expected reward $R(x)$ of samples $x$ from a pretrained diffusion model $p_{\theta}$, as shown in Eq. 1.
$\psi^{\star}=\underset{\psi\in\Psi}{\arg\max}\;\mathbb{E}_{x\sim p_{\theta}(x\mid\psi)}\bigl{[}R(x)\bigr{]}.$ (3)
Examples of control variables, $\psi$, include optimizing conditional input prompts [15, 12, 27, 10], manipulating cross-attention layers [11], and tuning latent noise vectors (noise optimization) to guide the diffusion trajectory [47, 42, 23, 45, 55, 26]. In this work, we focus exclusively on noise optimization ($\psi=z$), which is generalizable across diffusion models. Broadly, noise optimization methods fall into gradient-based optimization, or gradient-free optimization. We discuss each approach in turn.
Gradient-based methods refine the noise iteratively by leveraging the gradient with respect to the reward. Direct Optimization of Diffusion Latents, DOODL [47], and Direct Noise Optimization, DNO [42], are exemplary of these approaches. Both need to contend with the computational costs of backpropagation and maintaining sample quality by keeping the optimized noise on the Gaussian shell. DOODL optimizes the diffusion noise by computing the gradients with respect to a differentiable loss on generated images. They keep memory costs constant by leveraging invertible networks, and ensure sample quality by normalizing the optimized noise to have the norm of the original. Likewise, DNO computes the gradient through the reward function to optimize the noise, but can optimize non-differentiable rewards by using zero-th order optimization algorithms. The drawbacks of gradient-based methods are: they are not black-box; they are computationally expensive, especially when aligning multiple samples; and often require long optimization budgets (e.g., DNO requires $>100$ optimization steps, whereas our method achieves better results in $\sim$50 steps).
Gradient-free methods, on the other hand, explore the space of noise vectors or trajectories using only search or sampling methods. Ma et al. [26], employ three different strategies — random search, zero-order search (similar to hill climbing), and search-over-paths — to find samples that maximize their reward functions. Uehara et al. [45] provide a comprehensive overview of various inference-time algorithms, covering sequential Monte Carlo (SMC)-based guidance, value-based importance sampling, tree search, and classifier guidance. Li et al. propose DSearch [23], a dynamic beam search algorithm in order to search for noise in order to maximize some reward function. These techniques vary in their trade-offs, for example DSearch and the techniques outlined by Uehara et al. are not black-box and may have large computational costs. Meanwhile, although random and zero-order search are black-box and computationally efficient we show evolutionary methods outperform them (§ 4.2).
In this work, we investigate a different class of gradient-free optimization methods: evolutionary algorithms. We pursue evolutionary algorithms due to their: (1) ability to be used for black-box optimization; (2) overcoming certain computational and hardware limitations associated with existing gradient-based and complex search-based methods; (3) potential to effectively explore multi-modal search spaces. As such, despite being a black-box method, we outperform both classes of methods, while remaining computationally efficient.
B.2 Extended Alignment Objective Commentary
Here, we make additional comments upon Eq. 2 as presented in § 3. Recall that Eq. 2:
$\phi^{\star}=\arg\max_{\phi}\;\underbrace{\mathbb{E}_{\psi\sim q_{\phi}(\psi)}}_{\begin{subarray}{c}\text{search}\\ \text{distribution}\end{subarray}}\Bigl{[}\mathbb{E}_{x\sim p_{\theta}(x\mid\psi)}\bigl{[}R(x)\bigr{]}\Bigr{]}\,,$
cast inference-time alignment as search problem where we find the search distribution $q_{\phi}(\psi)$, that maximizes the expected reward. This objective is a generalization of Eq. 1 to natural evolutionary strategies [50]. Namely, we rather than searching for a single $\psi$ that maximizes the expected reward, we assume that $\psi$ is sampled from a parameterized search distribution, $q_{\phi}$, and we maximize the expected reward under this distribution. Note: $x$ is either sampled from the solution space $\psi$ in a stochastic manner or computed deterministically. In this context, $\theta$ is the parameterization of the sampler of the diffusion model, and the sampler is modeled either as $x\sim p_{\theta}(x|\psi)$ (stochastic) or $x=f_{\theta}(\psi)$ (deterministic).
B.3 Extended Characterization of Evolutionary Algorithm Components
Here, we provide an extended and in-depth overview of the mechanisms of genetic algorithms and natural evolutionary strategies.
Genetic Algorithms. Genetic algorithms maintain a population of candidate solutions, $\psi_{i}$, and iteratively improve them via selection, crossover, and mutation [8]. In the context of Eq. 2, the population is viewed as an empirical distribution. Thus the objective is to maximize the expected reward of a population of solutions. This objective is maximized iteratively and at each generation:
  1. 1.
    Evaluation: For each solution $\psi_{i}$ in the population, the fitness (reward) is evaluated — i.e., $R(x_{i})$ where $x_{i}\sim p_{\theta}(x|\psi_{i})$.
  2. 2.
    Selection: Parent vectors, $\psi_{i}$ are chosen based on fitness, e.g. via tournament or fitness-proportionate selection. In the context of alignment, selection will result in choosing better aligned solutions to use as parents for the next generation of solutions.
  3. 3.
    Crossover: Pairs of parents exchange noise component (or transformation parameters) to produce offspring. Correct choice of crossover can help ensure we meet (C1). For example, when using uniform crossover (swapping each coordinate independently), each child’s coordinate is drawn from one of two i.i.d. $\mathcal{N}(0,1)$ parents thus remaining with high-density shell (§ B.4).
  4. 4.
    Mutation: Offspring are perturbed with noise, often in an additive manner. E.g. $x=x+\epsilon,\;\text{where}\;\epsilon\sim\mathcal{N}(0,\sigma)$. $\sigma$ is a hyperparameter of the genetic algorithm.
Natural Evolutionary Strategies. Natural evolutionary strategies perform black-box optimization by adapting the parameterized search distribution $q_{\phi}$ over solutions [50]. Unlike GAs, $q_{\phi}$ is a parameterized search distribution (e.g., multivariate normal) and at each iteration its parameters are updated as follows:
  1. 1.
    Sampling: A population of solutions is sampled for the search distribution $\psi_{i}\sim q_{\phi}(\psi)$.
  2. 2.
    Evaluation: For each solution $\psi_{i}$ in the population, we evaluate the fitness as $R(x_{i})$, where $x_{i}\sim p_{\theta}(x|\psi_{i})$.
  3. 3.
    Update: Estimate the gradient $\nabla_{\phi}\mathbb{E}_{\psi\sim q_{\phi}(\psi)}\Bigl{[}\mathbb{E}_{x\sim p_{\theta}(x\mid\psi)}\bigl{[}R(x)\bigr{]}\Bigr{]}$ and update $\phi$ to increase expected reward along the natural gradient. Thus, we update the parameters of the search distribution $q_{\phi}$ so that future samples are better aligned.
B.4 Proof of Gaussian Marginal Preservation under Uniform Crossover
Purpose
In this section we show that uniform crossover maintains the solutions within the high-density Gaussian shell, which was alluded to in  § 3.3. This is relevant specifically to the EvoTorch implementation of the CoSyne algorithm, which uses a uniform distribution to decide crossover behavior.
Claim
Let $X=(X_{1},\dots,X_{d})$ and $Y=(Y_{1},\dots,Y_{d})$ be independent draws from $\mathcal{N}(0,I_{d})$. To perform coordinate-wise uniform crossover, for each $i$ we independently sample:
$B_{i}\sim\mathrm{Bernoulli}\bigl{(}p\bigr{)},\quad Z_{i}=\begin{cases}X_{i},&B_{i}=1,\\ Y_{i},&B_{i}=0.\end{cases}$
Then each coordinate $Z_{i}\sim\mathcal{N}(0,1)$ and thus $Z\sim\mathcal{N}(0,I_{d})$.
Proof.
Fix a coordinate $i$. For any real $z$, by the law of total probability conditioning on $B_{i}$:
$P(Z_{i}\leq z)$ $=P(Z_{i}\leq z\mid B_{i}=1)\,P(B_{i}=1)+P(Z_{i}\leq z\mid B_{i}=0)\,P(B_{i}=0)$
$=P(B_{i}=1)\,P(X_{i}\leq z)+P(B_{i}=0)\,P(Y_{i}\leq z)$
$=p\,\Phi(z)+(1-p)\,\Phi(z)$
$=\Phi(z)$
where $\Phi$ is the standard normal CDF and we used $X_{i},Y_{i}\sim\mathcal{N}(0,1)$. Thus the CDF of $Z_{i}$ matches that of $\mathcal{N}(0,1)$, so $Z_{i}\sim\mathcal{N}(0,1)$. Since each $B_{i}$ acts independently on its coordinate and all coordinates of $X$, $Y$ are independent, the $Z_{i}$ remain independent. Therefore the joint distribution of $Z$ is
$P(Z_{1}\leq z_{1},\dots,Z_{d}\leq z_{d})=\prod_{i=1}^{d}P(Z_{i}\leq z_{i})=\prod_{i=1}^{d}\Phi(z_{i}),$
which is the CDF of $\mathcal{N}(0,I_{d})$. Equivalently, $Z\sim\mathcal{N}(0,I_{d})$. ∎
Implication
As noted in § 3.3, uniform crossover will ensure that solutions are with the high-density shell of the gaussian thus maintaining sample quality.
Appendix C Details on Parameterization of Evolutionary Algorithms
Here, we review more details on how we initialized our search parameters. We also comment on consequences and insights related to choices of initialization.
GA Mutation Rate and Stability. When using GAs, the mutation operator introduces Gaussian noise to promote exploration. However, excessive perturbations can push solutions outside the high-density shell, leading to poor sample quality. To address this, we use small mutation step sizes, which help ensure that offspring remain within the high-density shell and preserve sample quality. Typically, our mutation size is $\sigma\approx 0.1$.
Evolutionary Strategy Initialization. Here, we briefly describe pitfalls with improper ES initialization, and how we addressed them in our work. Fig. 6 visually depicts this concept. We noted that ES are particularly sensitive to initialization — in the context of this work, this "initialization" is the initial search distribution $q_{\phi}(\psi)$.
One naive, but simple way to initialize $q_{\phi}(\psi)$ is to model it as a zero-mean ($\mu=0$), isotropic (covariance $\sigma=I$) multivariate Gaussian distribution. However, even after a few optimizations steps, the ES algorithm may update $\mu$ and $\sigma$ such that sampling from $q_{\phi}(\psi)$ no longer yields samples on the Gaussian shell. As a result, sample quality is degraded significantly. This is visually depicted in the top row of Fig. 6.
Figure 6: We depict a shortcoming with naive ES initialization, and illustrate our solution. Initializing the search distribution $q_{\phi}(\psi)$ incorrectly (zero-mean, isotropic multivariate Gaussian) can quickly lead to sampling outside the Gaussian shell, leading to poor sample quality. We encountered this issue in our early experiments. We addressed it by centering $q_{\phi}(\psi)$ on some point on the shell, and restricting its initial covariance $\sigma_{0}$ to a localized region, rather than encapsulating the entire shell.
In order to address this problem, we instead randomly choose an initial $\mu$ that exists on the Gaussian shell, and initialize the covariance $\sigma$ to be restricted to a localized region. This modification was effective at improving sample quality, because our samples were now more closely drawn from the surface of the Gaussian shell.
Appendix D Extended Evalaution Results
Here, we introduce additional quantiative and qualitative results.
D.1 Significance Testing
We verify whether our results are statistically significant compared to unguided, random sampling of noise by performing Wilcoxon’s signed rank test; which assumes that distribution of pair differences is symmetric. We compare the resulting $p$ value from Wilcoxon’s against the (standard) significance level $\alpha=0.05$. Entries in  Tab. 2 and  Tab. 1 are annotated with a dagger if they are statistically insignificant. Across both tables, only 8 out of 84 of the entries related to our method (CoSyne, SNES, PGPE) were insignificant.
D.2 Effect of longer optimization timelines for gradient-based methods
Figures 7, 8 and 9 depict the effect of longer optimization timelines on the best-sample, mean, and median rewards for the gradient-based method DNO, compared to gradient-free methods. Gradient-based require longer optimization timelines to reach the same reward as gradient-free methods.
Figure 7: DNO with Long Optimization Horizon: Best-sample reward per step measured (ImageReward) as measured on Open Image Preferences. Horizontal lines are the maximum reward achieved by the gradient-free methods in 50 steps.
Figure 8: DNO with Long Optimization Horizon: Mean reward per step measured (ImageReward) as measured on Open Image Preferences. Horizontal lines are the maximum (mean) reward achieved by the gradient-free methods in 50 steps.
Figure 9: DNO with Long Optimization Horizon: Median reward per step measured (ImageReward) as measured on Open Image Preferences. Horizontal lines are the maximum (median) reward achieved by the gradient-free methods in 50 steps.
D.3 Population Size Ablation
Figures 10, 11 and 12 depict the effect of increasing population size on the CoSyne, SNES, and PGPE algorithms. All algorithms benefit from larger population sizes, however PGPE appears to benefit the most.
(a) Best-sample reward per step.
(b) Mean reward per step.
(c) Median reward per step.
Figure 10: Genetic Algorithm: CoSyNE Reward statistics for the CosSyNE algorithm across optimization step.
(a) Best-sample reward per step.
(b) Mean reward per step.
(c) Median reward per step.
Figure 11: Evolutionary Strategy: SNES Reward statistics for the SNES algorithm across optimization step.
(a) Best-sample reward per step.
(b) Mean reward per step.
(c) Median reward per step.
Figure 12: Evolutionary Strategy: PGPE Reward statistics for the PGPE algorithm across optimization step.
D.4 Selection Pressure and Convergence
Fig. 13 shows the effect on increasing selection pressure (tournament size) on the median reward and standard deviation. We note that high selection pressure (larger tournament size) results in a rapid reduction in the reward standard deviation as shown in Fig. 13(b), this is in line with prior work characterizing the effect of tournament selection [30].
(a) Median reward per step.
(b) Reward standard deviation per step.
Figure 13: Genetic Algorithm: CoSyne Reward statistics per step, as measured on Open Image Preferences. Reward statistics are measured with increasing selection pressure, i.e., increasing tournament sizes of 2, 4, and 8. We fix the population size to 16.
D.5 Qualitative Results
Here we provide qualitative and visual results. We first address reward hacking behavior we witnessed with the PGPE algorithm optimizing for JPEG compressibility. We then present sets of images that illustrate the low sample diversity of genetic algorithms across models and datasets.
D.5.1 PGPE JPEG Reward-Hacking
Fig. 14 shows the proclivity of PGPE to "reward-hack" as noted in § 4.2. In many instances, PGPE was able to reduce the JPEG file size, but the resulting images were far too dissimilar or nonsensical compared with the intended image for a prompt. This behavior is reward-hacking, because the algorithm completely neglects the prompt, and instead produces nonsensical images to maximize reward.
Figure 14: PGPE Reward Hacking: Randomly selected images created using inference-time alignment with the PGPE algorithm (transformation search) to minimize the JPEG size (DrawBench). PGPE neglects the prompt, and naively attempts to minimize JPEG file size by producing monochromatic or simple images.
D.5.2 Stable Diffusion Qualitative Results
Here, we present sets of images from DrawBench and Open Image Preferences to illustrate qualitative results — Fig. 15, Fig. 16, Fig. 17, and Fig. 18. Specifically, we show how genetic algorithms such as CoSyne feature low sample diversity, even as optimization steps scale. This can also happen to ES methods (Fig. 18, however we noted it was less likely to occur with ES. Over longer optimization horizons, ES was able to maximize rewards better than GA while having higher diversity ( § 4.4).
Figure 15: Qualitative Sample Diversity: Randomly selected DrawBench prompts evaluated on StableDiffusion-1.5 with ImageReward. CoSyne was used to perform alignment. Across optimization steps, the diversity between samples quickly diminishes. We identify this as a consistent occurrence with genetic algorithms such as CoSyne.
Figure 16: Qualitative Sample Diversity: Randomly selected Open Image Preference prompts evaluated on StableDiffusion-1.5 with ImageReward. CoSyne was used to perform alignment. Note the low sample diversity.
Figure 17: Qualitative Sample Diversity: Randomly selected Open Image Preference prompts evaluated on StableDiffusion-3 with ImageReward. CoSyne was used to perform alignment. This low sample diversity result on StableDiffusion-3 was also noticed on StableDiffusion-1.5.
Figure 18: Qualitative Sample Diversity: Randomly selected Open Image Preference prompts evaluated on StableDiffusion-3 with ImageReward. SNES was used to perform alignment.
D.6 Extended DrawBench Results
Table 4: Extended DrawBench results with ImageReward.
(a) Population-Best Reward
Algorithm Soln. Space Mean $\pm$ Std Median Min Max
Random - $1.407\pm 0.519$ 1.583 -0.569 1.988
ZeroOrder - $1.455\pm 0.530$ 1.673 -0.408 2.002
DNO - $0.707\pm 0.926$ 0.910 -1.744 1.950
Cosyne Noise $1.614\pm 0.417$ 1.765 -0.093 2.002
Cosyne Rotation $1.532\pm 0.437$ 1.659 -0.259 2.001
SNES Noise $1.392\pm 0.523$ 1.535 -0.402 1.977
SNES Rotation $1.338\pm 0.561$ 1.482 -0.744 1.983
PGPE Noise $1.261\pm 0.646$ 1.419 -1.166 1.978
PGPE Rotation $1.276\pm 0.577$ 1.426 -0.641 1.981
(b) Median Reward
Algorithm Soln. Space Mean $\pm$ Std Median Min Max
Random - $0.383\pm 0.844$ 0.437 -1.717 1.877
ZeroOrder - $0.966\pm 0.733$ 1.093 -1.532 1.976
DNO - $0.707\pm 0.926$ 0.910 -1.744 1.950
Cosyne Noise $1.379\pm 0.573$ 1.565 -0.608 1.993
Cosyne Rotation $1.225\pm 0.633$ 1.420 -0.811 1.982
SNES Noise $0.788\pm 0.801$ 0.934 -1.387 1.944
SNES Rotation $0.942\pm 0.727$ 1.132 -1.541 1.954
PGPE Noise $0.921\pm 0.801$ 1.112 -1.681 1.951
PGPE Rotation $0.882\pm 0.731$ 1.003 -1.393 1.970
Table 5: Extended DrawBench results with HPSv2.
(a) Population-Best Reward
Algorithm Soln. Space Mean $\pm$ Std Median Min Max
Random - $0.301\pm 0.016$ 0.304 0.258 0.344
ZeroOrder - $0.304\pm 0.017$ 0.306 0.260 0.347
DNO - $0.286\pm 0.017$ 0.286 0.241 0.324
Cosyne Noise $0.310\pm 0.017$ 0.311 0.271 0.352
Cosyne Rotation $0.307\pm 0.017$ 0.308 0.267 0.347
SNES Noise $0.300\pm 0.016$ 0.301 0.260 0.346
SNES Rotation $0.300\pm 0.017$ 0.302 0.250 0.342
PGPE Noise $0.299\pm 0.017$ 0.301 0.256 0.342
PGPE Rotation $0.300\pm 0.017$ 0.301 0.251 0.338
(b) Median Reward
Algorithm Soln. Space Mean $\pm$ Std Median Min Max
Random - $0.282\pm 0.016$ 0.284 0.238 0.318
ZeroOrder - $0.290\pm 0.017$ 0.292 0.245 0.336
DNO - $0.286\pm 0.017$ 0.286 0.241 0.324
Cosyne Noise $0.300\pm 0.016$ 0.302 0.259 0.340
Cosyne Rotation $0.299\pm 0.017$ 0.300 0.262 0.340
SNES Noise $0.286\pm 0.016$ 0.288 0.247 0.324
SNES Rotation $0.290\pm 0.017$ 0.293 0.243 0.333
PGPE Noise $0.290\pm 0.017$ 0.291 0.252 0.335
PGPE Rotation $0.290\pm 0.017$ 0.294 0.243 0.329
Table 6: Extended DrawBench results with CLIP.
(a) Population-Best Reward
Algorithm Soln. Space Mean $\pm$ Std Median Min Max
Random - $37.139\pm 3.665$ 37.016 29.094 49.562
ZeroOrder - $37.590\pm 3.864$ 37.719 28.000 48.344
DNO - $26.237\pm 3.638$ 26.309 18.047 35.562
Cosyne Noise $38.791\pm 3.812$ 38.672 30.328 50.062
Cosyne Rotation $38.405\pm 3.795$ 37.938 30.125 50.188
SNES Noise $38.061\pm 3.751$ 37.875 29.094 49.562
SNES Rotation $37.407\pm 3.831$ 37.188 29.250 48.156
PGPE Noise $37.052\pm 3.607$ 36.656 29.250 47.844
PGPE Rotation $36.884\pm 3.763$ 36.609 28.562 48.875
(b) Median Reward
Algorithm Soln. Space Mean $\pm$ Std Median Min Max
Random - $32.867\pm 3.502$ 32.812 20.562 41.750
ZeroOrder - $34.545\pm 3.698$ 34.438 24.578 44.938
DNO - $26.237\pm 3.638$ 26.309 18.047 35.562
Cosyne Noise $36.455\pm 3.614$ 36.141 28.219 46.938
Cosyne Rotation $36.393\pm 3.712$ 36.266 26.266 47.250
SNES Noise $35.623\pm 3.563$ 35.328 24.547 44.969
SNES Rotation $34.699\pm 3.662$ 34.438 26.594 45.812
PGPE Noise $34.397\pm 3.455$ 34.031 25.859 43.469
PGPE Rotation $34.772\pm 3.602$ 34.516 27.172 44.750
Table 7: Extended DrawBench results with JPEG File Size. Entries are file sizes listed in kilobytes (kB).
(a) Population-Best Reward
Algorithm Soln. Space Mean $\pm$ Std Median Min Max
Random - $66.92\pm 17.82$ 64.23 26.55 136.58
ZeroOrder - $53.35\pm 18.42$ 51.47 15.89 116.90
DNO - $94.64\pm 22.78$ 91.78 47.11 174.45
Cosyne Noise $39.52\pm 15.88$ 36.79 13.78 79.34
Cosyne Rotation $38.73\pm 16.25$ 36.82 9.96 108.24
SNES Noise $71.35\pm 21.36$ 67.54 28.47 157.80
SNES Rotation $57.78\pm 22.40$ 54.63 11.54 174.64
PGPE Noise $88.29\pm 23.68$ 84.50 31.79 160.11
PGPE Rotation $18.29\pm 12.37$ 14.76 5.00 91.34
(b) Median Reward
Algorithm Soln. Space Mean $\pm$ Std Median Min Max
Random - $105.43\pm 20.66$ 103.54 60.79 203.58
ZeroOrder - $62.76\pm 20.65$ 61.21 18.32 158.85
DNO - $94.64\pm 22.78$ 91.78 47.11 174.45
Cosyne Noise $44.76\pm 16.40$ 42.42 15.42 87.25
Cosyne Rotation $42.28\pm 16.63$ 40.86 11.91 111.33
SNES Noise $77.98\pm 22.33$ 73.70 32.36 168.50
SNES Rotation $66.10\pm 23.36$ 62.25 16.72 185.45
PGPE Noise $97.02\pm 24.02$ 93.35 53.09 173.79
PGPE Rotation $20.18\pm 13.10$ 16.60 5.38 94.78
Here, we include additional results on the DrawBench dataset. Specifically, we include additional statistics (min, max, median, mean, standard) for each reward function. Each subtable for a particular reward function displays the population-best rewards and median rewards, which were measured across all prompts of the DrawBench dataset. We provide this data in Tab. 4, Tab. 5, Tab. 6, and Tab. 7.

### v2

@@
-Inference Time Alignment of Diffusion Models with Evolutionary Algorithms
-Inference Time Alignment of Diffusion Models with Evolutionary Algorithms
-Purvish Jajal*
-Purdue University
-West Lafayette, IN 47907
-pjajal@purdue.edu
-&Nick John Eliopoulos*
-Purdue University
-West Lafayette, IN 47907
-neliopou@purdue.edu
-&Benjamin Shiue-Hal Chou
-Purdue University
-West Lafayette, IN 47907
-chou150@purdue.edu
-&George K. Thiruvathukal
-Loyola University Chicago
-Chicago, IL 60660
-gkt@cs.luc.edu
-&James C. Davis
-Purdue University
-West Lafayette, IN 47907
-davisjam@purdue.edu
-&Yung-Hsiang Lu
-Purdue University
-West Lafayette, IN 47907
-yunglu@purdue.edu
-Evolutionary Algorithms for Inference-Time Alignment of Diffusion Models
-Purvish Jajal*
-Purdue University
-West Lafayette, IN 47907
-pjajal@purdue.edu
-&Nick John Eliopoulos*
-Purdue University
-West Lafayette, IN 47907
-neliopou@purdue.edu
-&Benjamin Shiue-Hal Chou
-Purdue University
-West Lafayette, IN 47907
-chou150@purdue.edu
-&George K. Thiruvathukal
-Loyola University Chicago
-Chicago, IL 60660
-gkt@cs.luc.edu
-&James C. Davis
-Purdue University
-West Lafayette, IN 47907
-davisjam@purdue.edu
-&Yung-Hsiang Lu
-Purdue University
-West Lafayette, IN 47907
-yunglu@purdue.edu
-Evolutionary Algorithms for Black-Box Inference-Time Alignment of Diffusion Models
-Purvish Jajal*
-Purdue University
-West Lafayette, IN 47907
-pjajal@purdue.edu
-&Nick John Eliopoulos*
-Purdue University
-West Lafayette, IN 47907
-neliopou@purdue.edu
-&Benjamin Shiue-Hal Chou
-Purdue University
-West Lafayette, IN 47907
-chou150@purdue.edu
-&George K. Thiruvathukal
-Loyola University Chicago
-Chicago, IL 60660
-gkt@cs.luc.edu
-&James C. Davis
-Purdue University
-West Lafayette, IN 47907
-davisjam@purdue.edu
-&Yung-Hsiang Lu
-Purdue University
-West Lafayette, IN 47907
-yunglu@purdue.edu
-Inference-Time Alignment of Diffusion Models with Evolutionary Algorithms
-Purvish Jajal*
-Purdue University
-West Lafayette, IN 47907
-pjajal@purdue.edu
-&Nick John Eliopoulos*
-Purdue University
-West Lafayette, IN 47907
-neliopou@purdue.edu
-&Benjamin Shiue-Hal Chou
-Purdue University
-West Lafayette, IN 47907
-chou150@purdue.edu
-&George K. Thiruvathukal
-Loyola University Chicago
-Chicago, IL 60660
-gkt@cs.luc.edu
-&James C. Davis
-Purdue University
-West Lafayette, IN 47907
-davisjam@purdue.edu
-&Yung-Hsiang Lu
-Purdue University
-West Lafayette, IN 47907
-yunglu@purdue.edu
+Inference-Time Alignment of Diffusion Models via Evolutionary Algorithms
 Abstract
-Diffusion models are state-of-the-art generative models in various domains, yet their samples often fail to satisfy downstream objectives such as safety constraints or domain-specific validity.
-Existing techniques for alignment require gradients, internal model access, or large computational budgets.
-We introduce an inference-time alignment framework based on evolutionary algorithms.
+Diffusion models are state-of-the-art generative models, yet their samples often fail to satisfy application objectives such as safety constraints or domain-specific validity.
+Existing techniques for alignment require gradients, internal model access, or large computational budgets — resulting in high compute demands, or lack of support for certain objectives.
+In response, we introduce an inference‑time alignment framework based on evolutionary algorithms.
 We treat diffusion models as black-boxes and search their latent space to maximize alignment objectives.
-Our method enables efficient inference-time alignment for both differentiable and non-differentiable alignment objectives across a range of diffusion models.
-On the DrawBench and Open Image Preferences benchmark, our EA methods outperform state-of-the-art gradient-based and gradient-free inference-time methods.
-In terms of memory consumption, we require $55\%$ to $76\%$ lower GPU memory than gradient-based methods.
-In terms of running-time, we are $72\%$ to $80\%$ faster than gradient-based methods.
-We achieve higher alignment scores over 50 optimization steps on Open Image Preferences than gradient-based and gradient-free methods.
-footnotetext: * Purvish Jajal and Nick John Eliopoulos contributed equally to this work.
+Given equal or less running time, our method achieves 3-35% higher ImageReward scores than gradient-free and gradient-based methods.
+On the Open Image Preferences dataset, our methods achieve competitive results across four popular alignment objectives.
+In terms of computational efficiency, we require 55% to 76% less GPU memory and are 72% to 80% faster than gradient-based methods.
+footnotetext: * P. Jajal and N.J. Eliopoulos contributed equally to this work.
 1 Introduction
-Figure 1: Samples generated by our method on Stable Diffusion-3 across three different prompts under various alignment objectives (brightness, saturation, file size).
+Figure 1: Samples generated by our method on Stable Diffusion-3.
 Each row shows the progression over optimization steps $i$, with the corresponding metric values displayed in the top-right corner.
-The noise in each generation is optimized via the CoSyNE algorithm [13] over 14 optimization steps.
-Arrows ($\uparrow$/$\downarrow$) indicate whether the metric is being maximized or minimized.
-Diffusion models [40, 34, 9] have emerged as state-of-the-art generative models for synthesizing high-quality images, video, and audio [3, 16, 18, 25].
-However, the outputs of diffusion models often fail to meet downstream objectives, such as adhering to user preferences [53, 51, 37], safety constraints [28], or domain-specific validity [14, 49].
-This problem is referred to as alignment—ensuring that generated samples satisfy external objectives beyond the original maximum-likelihood objective [24, 48, 2, 21, 5, 31, 4].
-Alignment is an important concern for many practical applications such as image generation [53, 51, 37] and drug discovery [14, 49].
-Methods for diffusion model alignment can be categorized into two types: fine-tuning-based alignment and inference-time alignment.
-Fine-tuning-based alignment typically relies on supervised fine-tuning or reinforcement learning to adjust model parameters [24].
-In contrast, inference-time alignment modifies the sampling process during generation to maximize a reward function (i.e., alignment objective) [45].
-These methods aim to be computationally efficient and support arbitrary rewards while maintaining sample quality.
-However, existing approaches have several drawbacks: they are often not black-box, requiring access to gradients or internal model states; they cannot efficiently generate many aligned samples; and they are often not sample-efficient.
-As a result, they are difficult to integrate with a variety of models, sampling strategies, and reward functions; generating many aligned samples becomes cost-prohibitive; and obtaining aligned samples may take too long.
-In this work, we demonstrate that evolutionary algorithms (EAs)—specifically genetic algorithms (GA) and evolutionary strategies (ES)—can be used to align diffusion models.
-As a class of algorithms, EAs have desirable properties: they are black-box, gradient-free optimizers, that can be highly parallelized [39, 36].
-By interpreting the diffusion models as black-boxes and searching over their latent space by using EAs we can optimize the alignment objective.
-Our approach enables efficient alignment that supports both differentiable and non-differentiable rewards, and is compatible with various diffusion models.
-Fig.
-1 depicts our method applied to various alignment objectives.
-To quantify alignment quality, GPU-memory footprint, and runtime efficiency, we evaluate our inference-time alignment method on two prompt collections [35, 6], four reward functions [53, 51, 32, 4], and three diffusion models [34, 9, 3].
-On both prompt collections, our method outperforms existing gradient-free and gradient-based inference-time alignment approaches.
-In terms of computational costs, we exhibit $55\%$ to $76\%$ reduction in GPU memory than gradient-based methods, while only consuming up to 1.5×1.5$\times$1.5 × more memory than gradient-free methods.
-When producing the same number of aligned samples, evolutionary methods achieve 3.5×3.5$\times$3.5 × to $\times5$ lower runtime per optimization step compared to gradient-based methods, and only $\times1$ to 1.6×1.6$\times$1.6 × higher running time per step than gradient-free methods.
-Moreover, evolutionary methods reach higher alignment scores than gradient-free and gradient-based methods in 10 optimization steps.
-In sum, we contribute:
-  • We propose a novel inference-time alignment approach using evolutionary algorithms (EAs).
-Specifically, we introduce two black-box methods for aligning the outputs of diffusion models: (1) optimizing noise, and (2) optimizing transformations applied to the noise.
-We elaborate upon the requisite choices for utilizing EAs for inference-time alignment, such as search algorithms, initialization, operators, and computational considerations.
-  • We demonstrate the effectiveness of two widely used classes of evolutionary algorithms on the diffusion model alignment task: genetic algorithms and evolutionary strategies.
-Our evaluation spans a diverse set of reward functions and diffusion models.
-2 Related Work
-In this section, we review prior work on diffusion model alignment (§ 2.1) and evolutionary algorithms (§ 2.2). An extended treatment of related works can be found in § B.1.
+The noise in each generation is optimized via our evolutionary-based approach, which uses the CoSyNE algorithm [gomez2008accelerated] over 14 optimization steps.
+Arrows ($\uparrow$$\downarrow$) indicate whether the metric is being maximized or minimized.
+Diffusion models [sohl-dickstein_deep_2015, rombach2022high, esser2024scaling] are state-of-the-art generative models for synthesizing high-quality images, video, and audio [chen2023pixart, ho2022video_diffuion, kong2020diffwave, liu2024latent].
+However, diffusion model outputs often fail to meet downstream objectives, such as user preferences [xu2023imagereward, hpsv2wu2023human, laionAestheticDataset], safety constraints [miao2024t2vsafetybench], or domain-specific validity [gu2024aligning, wang2024fine].
+This problem is referred to as alignment—ensuring generated samples satisfy objectives beyond the original maximum-likelihood objective [liu2024alignmentdiffusionmodelsfundamentals, wallace2024diffusion, black2023training, lee2023aligning, dai2023emu, prabhudesai2023aligning, clark2023directly].
+We reiterate the desirable qualities of a diffusion alignment method [uehara2025inference, tang2024inference_dno]: (1) black-box, meaning it applies on a wide range of models because access to gradients or internal model states is not required; (2) support for arbitrary alignment objectives, (3) maintaining sample-efficiency, meaning an objective can be optimized with few samples.
+Most prior works capture only some of these properties [uehara2025inference].
+The works that satisfy all these properties (Best-of-N, Zero-Order) are often used as simple baselines, and we find that these baselines can be competitive with white-box methods.
+This suggests that alternate, black-box alignment formulations are worth investigating.
+In this work, we ask whether black-box evolutionary algorithms (EAs) can be used for diffusion model alignment.
+EAs should satisfy [eiben2015introduction] the aforementioned criteria, though they have not been applied to the diffusion alignment task.
+We treat the diffusion model as a black box, and evolve its latent noise vector with EAs to maximize alignment objectives.
+This approach enables efficient alignment supporting both differentiable and non-differentiable rewards, and is compatible with various diffusion models.
+Fig.˜1 depicts our method applied to various alignment objectives.
+We evaluate our method in terms of its ability to achieve higher rewards than other methods under short-term optimization and equal-compute regimes.
+Our method is evaluated with two prompt datasets [drawbench2022, data_is_better_together_open_image_preferences_v1_flux_dev_lora_2024], four objective functions [xu2023imagereward, hpsv2wu2023human, clip2021, clark2023directly] (differentiable and non-differentiable), and on Stable Diffusion 1.5 [rombach2022high] as in prior work.
+Our method outperforms existing gradient-free and gradient-based inference-time alignment approaches on both prompt datasets.
+We achieve better sample-efficiency, rewards, and memory efficiency over short optimization horizons and under equal-compute regimes.
+Given similar or less running time, our method achieves 3-35% higher ImageReward scores than gradient-free and gradient-based methods.
+On the Open Image Preferences dataset, our methods achieve the highest rewards across four popular alignment objectives.
+In terms of computational efficiency, we require 55% to 76% lower GPU memory and are 72% to 80% faster than gradient-based methods, while consuming up to 1.51.5x more memory than gradient-free methods.
+In sum, our contributions are:
+  • We propose an inference-time alignment approach using evolutionary algorithms (EAs).
+Specifically, we introduce two black-box methods for aligning the outputs of diffusion models: (1) directly optimizing latent noise vectors, and (2) optimizing transformations applied to the noise vector.
+We elaborate on the considerations in using EAs for inference-time alignment, such as search algorithms, initialization, operators, and computational considerations.
+  • We evaluate two representative families of EAs (genetic algorithms and evolutionary strategies) for diffusion model alignment.
+Our evaluation shows generalizability over four popular reward functions, demonstrating EAs are both sample- and computationally-efficient.
+2 Background and Related Work
+In this section, we review prior work on diffusion model alignment (Sec.˜2.1) and evolutionary algorithms (Sec.˜2.2). Sec.˜B.1 gives an extended treatment of related works.
 2.1 Diffusion Model Alignment
-Diffusion models use a reverse diffusion process to convert some latent noise distribution into a data distribution, such as images [40, 16].
-The reverse diffusion process iteratively denoises the initial latent noise $z_{T}$ over some number of steps $T$ to yield a sample, $z_{0}$.
-Though diffusion models are capable of modeling complex data distributions, they often fail to produce samples that meet some downstream objective.
+Diffusion models use a reverse diffusion process to convert a latent noise distribution into a data distribution, such as images [sohl-dickstein_deep_2015, ho2022video_diffuion].
+The reverse diffusion process (sampling) iteratively denoises the initial latent noise $z_{T}$ over some number of steps $T$ to yield a sample, $z_{0}$.
+This denoising is guided by conditioning on some variable, usually a text-based prompt.
+Despite this, they may fail to produce samples that meet some downstream objective.
 Diffusion model alignment methods adjust diffusion models such that the resulting samples better meet an objective beyond the model’s original maximum likelihood criterion.
-These objectives commonly focus on producing images that reflect human aesthetic preferences [17, 51, 52, 53], or other metrics such as compressibility [2].
-Alignment methods can generally be grouped into two main categories: fine-tuning–based methods and inference-time methods.
-Fine-tuning-based alignment methods involve adjusting the diffusion model’s parameters so that generated samples better match alignment objectives [21, 5, 31, 4, 48, 2].
+Examples include aesthetics [kirstain2023pick, hpsv2wu2023human, hpswu2023human, xu2023imagereward], or compressibility [black2023training].
+Alignment methods can be grouped into two broad categories: fine-tuning-based methods and inference-time methods [uehara2025inference].
+Fine-tuning-based alignment methods involve adjusting the diffusion model’s parameters so that generated samples better match alignment objectives [lee2023aligning, dai2023emu, prabhudesai2023aligning, clark2023directly, wallace2024diffusion, black2023training].
 Fine-tuning based methods, although powerful, require retraining and thus all the associated costs.
 We follow the alternative approach, inference-time methods.
-Rather than altering model parameters, these techniques adjust sampling or conditioning to ensure samples meet alignment objectives.
-Formally, they seek the control variable $\psi$ that maximizes the expected reward $R(x)$ of samples $x$ from a pretrained diffusion model $p_{\theta}$, as shown in Eq.
-1.
-$\psi^{\star}=\underset{\psi\in\Psi}{\arg\max}\;\mathbb{E}_{x\sim p_{\theta}(x\mid\psi)}\bigl{[}R(x)\bigr{]}.$ (1)
-Examples of control variables, $\psi$, include optimizing conditional input prompts [15, 12, 27, 10], manipulating cross-attention layers [11], and tuning latent noise vectors (noise optimization) to guide the diffusion trajectory [47, 42, 23, 45, 55, 26].
-In this work, we focus exclusively on noise optimization ($\psi=z$), which is the most general method.
-Broadly, noise optimization methods fall into gradient-based optimization, or gradient-free optimization.
-We discuss each approach in turn.
+Rather than altering model parameters, these techniques adjust the diffusion sampling or conditioning to ensure samples meet alignment objectives.
+Formally, they seek the control variable $\psi$ that maximizes the expected reward $R(x)$ of samples $x$ from a pretrained diffusion model $p_{\theta}$, as shown in Eq.˜1.
+$\psi^{\star}=\underset{\psi\in\Psi}{\arg\max}\;\mathbb{E}_{x\sim p_{\theta}(x\mid\psi)}\bigl[R(x)\bigr]$ (1)
+Examples of control variables $\psi$ include optimizing conditional input prompts [hao2023optimizing, gal2022image, manas2024improving, fei2023gradient], manipulating cross-attention layers [feng2022training], and tuning latent noise vectors (noise optimization) to guide the diffusion trajectory [wallace2023end, tang2024inference_dno, li2025dynamic, uehara2025inference, zhou2024golden, ma2025inference_time_scaling].
+The most general control variable is noise optimization ($\psi=z$), which has two main categories: gradient-based and gradient-free methods.
 Gradient-based methods refine the noise iteratively by leveraging the gradient with respect to the reward.
-Direct Optimization of Diffusion Latents (DOODL) [47] and Direct Noise Optimization (DNO) [42] are exemplary of these approaches.
-Both need to contend with the computational costs of backpropagation and maintaining sample quality by keeping the optimized noise on the Gaussian shell—noise sampled from high-dimensional Gaussians concentrate on the surface of a spherical shell [46].
-The drawbacks of gradient-based methods are: they are not black-box; they are computationally expensive, when aligning multiple samples; and require long optimization budgets.
-Gradient-free methods, on the other hand, explore the space of noise vectors or trajectories using only search or sampling methods.
-Uehara et al.
-[45] provide a comprehensive overview of various inference-time algorithms, largely covering sampling-based methods.
-Ma et al.
-[26] employ three different sample search strategies — random search, zero-order search (similar to hill climbing), and search-over-paths.
+Recent examples include DOODL [wallace2023end] and Direct Noise Optimization (DNO) [tang2024inference_dno].
+These methods can incur large runtime and memory costs due to backpropagation, but can achieve high rewards over long optimization horizons.
+Gradient-free methods, on the other hand, explore the space of noise vectors or trajectories using search or sampling methods.
+Uehara et al[uehara2025inference] review inference-time algorithms, covering sampling-based and search-based methods.
+Ma et al[ma2025inference_time_scaling] employ three different search strategies — Best-of-N, Zero-Order, and Search over Paths.
 Li et al.
-propose DSearch [23], a dynamic beam search algorithm.
-These techniques vary in their trade-offs.
-DSearch and the techniques outlined by Uehara et al.
-are not black-box and may have large computational costs.
-Although random and zero-order search are also black-box and computationally efficient, evolutionary methods outperform them (§ 4.2).
+propose DSearch [li2025dynamic] and SVDD [li_svdd_2024], which are beam search algorithms.
+Singhal et al.
+propose Fk Steering [singhal_fk_steering_2025] (FKS), a sampling-based alignment method.
+DSearch, SVDD, and FKS methods are white-box, because they modify and rely on the diffusion denoising process.
+Although Best-of-N and Zero-Order search are black-box, we show that evolutionary methods outperform them in alignment scores and sample-efficiency (Secs.˜4.2 and 4.5).
+Figure 2: Mapping between Eq.˜2 and Algorithm˜1 search over $z_{T}$ directly, or an affine transform of $z_{T}$.
+We depict connections between Eq.˜2 and our method via color-coding.
+We perform alignment on human preferences (HPSv2, ImageReward), JPEG size, and CLIP scores.
 2.2 Evolutionary Algorithms
-Evolutionary algorithms (EAs) are a class of biologically inspired methods that can be used to solve a variety of black-box optimization problems [8, 39].
-They iteratively select, recombine, and mutate populations of solutions to improve their fitness (quality of the solution).
-Among the various EAs, genetic algorithms (GAs) [8, 39] and natural evolutionary strategies (ES) [50] are most pertinent to this work.
-Genetic algorithms and natural evolutionary strategies have been widely applied in deep learning and deep reinforcement learning.
-In deep learning, genetic algorithms have facilitated neural architecture search and hyperparameter tuning [33, 29, 54].
-In deep reinforcement learning, both genetic algorithms and natural evolutionary strategies have been shown to be competitive with back-propagation based methods such as policy gradients [41, 36].
-GA and ES are also used in image generation [22, 38, 43, 10] and artificial life simulations [20].
-However, to our knowledge, we are the first work to apply EAs for black-box diffusion model alignment.
-3 Inference Time Alignment via Evolutionary Algorithms
-We cast inference-time alignment as a black-box search problem and apply an evolutionary search framework.
-For a pretrained diffusion model $p_{\theta}$ and reward function $R(x)$, we aim to find parameters $\phi$ of the search distribution $q_{\phi}(\psi)$ that maximizes the expected reward in Eq.
-2.
+Evolutionary algorithms (EAs) are a class of biologically inspired methods that can be used to solve a variety of black-box optimization problems [eiben2015introduction, simon2013evolutionary].
+They iteratively select, recombine, and mutate populations of solutions to improve their fitness (solution quality).
+Among the various EAs, genetic algorithms (GAs) [eiben2015introduction, simon2013evolutionary] and natural evolutionary strategies (ES) [wierstra2014natural] are well-studied and pertinent to this work.
+More details on GA and ES are in Appendix˜B.
+GA and ES have been widely applied in deep learning and deep reinforcement learning.
+In deep learning, genetic algorithms have facilitated neural architecture search and hyperparameter tuning [miikkulainen2024evolving, young2015optimizing].
+In deep reinforcement learning, both genetic algorithms and natural evolutionary strategies have been shown to be competitive with back-propagation-based methods such as policy gradients [such2017deep, salimans2017evolution].
+GA and ES are also used in image generation [lehman2011evolving, secretan2008picbreeder, tian2022modern, fei2023gradient] and artificial life simulations [kumar2024automating].
+We are the first to apply and parameterize EAs for black-box diffusion model alignment.
+3 Inference-Time Alignment via Evolutionary Algorithms
+We cast inference‐time alignment as a black-box search problem and apply an evolutionary search framework. Fig.˜2 visually depicts our method, which we now explain.
+Given a pretrained diffusion model $p_{\theta}$ and reward function $R(x)$ corresponding to some alignment objective, we aim to find parameters $\phi$ of the search distribution $q_{\phi}(\psi)$ that maximizes the expected reward in Eq.˜2.
 $\phi^{\star}=\arg\max_{\phi}\;\underbrace{\mathbb{E}_{\psi\sim q_{\phi}(\psi)}}_{\begin{subarray}{c}\text{search}\\ \text{distribution}\end{subarray}}\Bigl[\mathbb{E}_{x\sim p_{\theta}(x\mid\psi)}\bigl[R(x)\bigr]\Bigr]\,.$ (2)
-For generality, we write $x\sim p_{\theta}(x\mid\psi)$ due to many diffusion samplers being stochastic.
-In practice, we treat this as a deterministic map from control variable to sample, i.e., $f_{\theta}:\psi\mapsto x$.
-See § B.2 for further details.
-In § 3.1, we review design goals and requirements for our method.
-In § 3.2 we define the solution spaces ($\psi$), over both noise and noise-transformations.
-In § 3.3 we describe how genetic algorithms and natural evolutionary strategies define the search distribution $q_{\phi}(\psi)$ and how they optimize $\phi$ in Eq.
-2.
-Finally, § 3.4 presents our implementation and computational considerations.
-3.1 Design Criteria
-Prior works [45, 42] have outlined a set of Criteria that characterize effective inference-time alignment approaches: (C1) maintaining sample quality; (C2) computationally efficient; and (C3) supporting non-differentiable reward functions.
-In this work, we advocate a fourth criterion: (C4) alignment methods should strive to be black-box.
-Black-box inference-time alignment can be easily integrated with a variety of models, sampling strategies, and reward functions.
-For example, a black-box alignment method could be applied to a quantized diffusion model for fast inference, and could simultaneously support arbitrary reward functions.
-Figure 2: Illustration of our evolutionary search approach to noise vector optimization. Here, we depict the connection between terms of Eq. 2 and components of our method via color-coding.
+[…]
+In Sec.˜3.1 we define the solution spaces ($\psi$), over both noise and noise‐transformations.
+In Sec.˜3.2 we describe how genetic algorithms and natural evolutionary strategies define the search distribution $q_{\phi}(\psi)$ and how they optimize $\phi$ in Eq.˜2.
+Sec.˜3.3 discusses implementation considerations.
 3.1 Defining the Solution Space
-Before applying evolutionary algorithms, we must define the solution space, i.e., what we will search over, $\psi$.
-For aligning diffusion models, we explore two options: (1) directly searching over noise vectors, and (2) searching over transformations of an initial noise vector.
-In both cases, our overarching goal is to identify a solution that maximizes Eq.
-2.
+Before applying evolutionary algorithms, we must define the solution space $\psi$, i.e.
+what we will search over.
+For the diffusion alignment problem, we explore two options: (1) directly searching over noise vectors, and (2) searching over transformations of an initial noise vector.
+In both cases, our overarching goal is to identify a solution that maximizes Eq.˜2.
 Searching over Noise.
-The simplest solution space we can define is over the initial noise, i.e., $\psi=z_{T}^{\prime}$ in Eq.
-2.
+The simplest solution space we can define is over the initial noise, i.e.
+$\psi=z_{T}^{\prime}$ in Eq.˜2, as depicted in Algorithm˜1.
 Concretely, we define the search variable $\psi$ to be the initial noise vector, $z_{T}^{\prime}\sim q_{\phi}(z_{T}^{\prime})$, and optimize $\phi$ to maximize the expected reward.
-This direct search is straightforward, but hinges on proper initialization and careful tuning of mutation; without these, evolved solutions can drift into low-density regions of the latent space, leading to poor sample quality.
+This direct search is straightforward, but hinges on careful parameterization; otherwise solutions can drift into low-density regions of the latent space, leading to poor sample quality (Sec.˜3.3).
+Algorithm 1 Alignment via Direct Noise Search
+1:Pretrained diffusion model $p_{\theta}$, reward $R(\cdot)$, iterations $T$, population size $M$
+2:Search distribution $q_{\phi}(\psi)$: (GA) population $\{\psi_{i}\}_{i=1}^{M}$ or (ES) distribution $\mathcal{N}(\mu,\Sigma)$
+3:Initialization: (GA) $\psi_{i}\sim\mathcal{N}(0,I)$ or (ES) $\mu=z_{0}\sim\mathcal{N}(0,I)$, and $\Sigma=\sigma_{0}I$ with a small $\sigma_{0}$
+4:for $t=1,\dots,T$ do
+5:  Sample candidates $\{\psi_{i}\sim q_{\phi}\}_{i=1}^{M}$ $\triangleright$ initial noises $z_{T}^{\prime}$
+6:  Generate samples $\{x_{i}\}_{i=1}^{M}$ with $x_{i}=f_{\theta}(\psi_{i})$
+7:  Compute rewards $\{r_{i}\}_{i=1}^{M}$ with $r_{i}=R(x_{i})$
+8:  $\phi\leftarrow\text{EAUpdate}(\phi,\{\psi_{i}\}_{i=1}^{M},\{r_{i}\}_{i=1}^{M})$
+9:end for
+10:return final $q_{\phi}$ or best candidate
+NB: EAUpdate (line 7) depends on the specific evolutionary algorithm.
 Searching for Noise Transformations.
-Instead of searching over noise vectors directly, we can define an affine transformation of an initial noise vector ($z_{T}$) and search over its parameters.
-We define the transformed noise as $z_{T}^{\prime}=f(z_{T})=Az_{T}+b$, where $A\in\mathbb{R}^{d\times d}$, $z_{T},b\in\mathbb{R}^{d}$.
-Thus, in Eq.
-2 we set $\psi=[A,b]$ with $[A,b]\sim q_{\phi}([A,b])$ and optimize $\phi$ to maximize the expected reward in Eq.
-2 under $x\sim p_{\theta}(x\mid Az_{T}+b)$.
-[…]
-Choosing between noise and noise transformation search.
-The choice between noise and noise-transformation search hinges on the type of regularization each imposes and their relative computational costs.
-Noise search, while being the most direct and straightforward way to define our search space, does not confine solutions to the high-density shell of the Gaussian.
-Consequently, if the search algorithms (§ 3.3) fail to ensure shell-confinement, sample quality degrades.
-In contrast, searching over noise transformations will guarantee (given $b=0$, see § 3.2) residence within the high-density shell regardless of the choice of evolutionary algorithm.
-However, noise-transformation search incurs extra cost for factorization and linear transforms; we describe mitigation strategies in § 3.4.
+Alternatively, to ensure residence in the valid latent space, we can instead search for affine transformations of the initial noise vector ($z_{T}$).
+This requires only slight modification to Algorithm˜1 (Sec.˜B.4).
+We define the transformed noise as $z_{T}^{\prime}=f(z_{T})=Az_{T}+b$, where $A\in\mathbb{R}^{d\times d}$, and $z_{T},b\in\mathbb{R}^{d}$.
+Thus, in Eq.˜2 we set $\psi=[A,b]$ with $A,b]\sim q_{\phi}([A,b])$ and optimize $\phi$ to maximize the expected reward in Eq.˜2 under $x\sim p_{\theta}(x\mid Az_{T}+b)$.
+In general, $z_{T}^{\prime}$ may not reside in the high-density shell of $\mathcal{N}(0,I)$, leading to poor sample quality.
+To keep $z_{T}^{\prime}$ within the high-density shell, we can ensure $A$ is orthonormal.
+If $A$ is orthonormal and $b=0$, $z_{T}^{\prime}$ stays within the high-density shell of $\mathcal{N}(0,I)$ by the rotational invariance property of the Gaussian [gaussian_annulus_lecture_2017].
+In this work, we use the QR decomposition to extract the orthonormal component $Q(A)$, and apply $Q(A)$ to $z_{T}$.
+Importantly, we perform QR decomposition only on the channel dimension of $z_{T}$ to avoid large compute costs.
+Direct Noise vs.
+Noise Transform Search.
+Direct noise search, while being the straightforward way to define our search space, does not confine solutions to the high-density shell of the Gaussian.
+If the search algorithm fails to ensure shell-confinement, sample quality degrades.
+In contrast, searching over orthonormal noise transformations will guarantee residence within the high-density shell given $b=0$, as discussed above.
 3.2 Searching the Solution Space
-Given our solution space ($\psi$) we now turn to how we define the search distribution and find its parameters, i.e., $q_{\phi}(\psi)$.
-Our requirements (C1C4) call for black-box, gradient-free optimizers that preserve sample quality and support arbitrary rewards.
-Among the large family of optimization approaches [7, 19, 8], evolutionary algorithms (EAs) satisfy these criteria and have a history of efficacy in machine learning contexts (§ 2.2).
-We consider two families of EA: genetic algorithms (GAs) and natural evolutionary strategies (ES).
-For each, we first review their core mechanics, then describe how these mechanics map to solving the diffusion alignment task.
-Their specific optimization procedures are outlined in § B.3.
+Given the solution space ($\psi$), we now turn to how we define the search distribution and its parameters $q_{\phi}(\psi)$.
+We consider two major families of evolutionary algorithms (EAs) that have found prior use in DL: genetic algorithms (GAs) and natural evolutionary strategies (ES) [kramer2017genetic, eiben2015introduction].
+GA and ES should satisfy the criteria laid out in Sec.˜1, namely: black-box, sample-efficient, and supporting arbitrary rewards [kramer2017genetic, eiben2015introduction].
+We first review their core mechanics, then map them onto the diffusion alignment task.
+Their optimization procedures are outlined in Sec.˜B.3.
 Genetic Algorithms (GAs).
-Genetic algorithms maintain a population of candidate solutions, $\psi_{i}$, and iteratively improve them via selection, crossover, and mutation [8].
-In Eq.
-2, the search distribution is an empirical distribution over the population of solutions.
-In the context of alignment, GAs satisfy the criteria in § 3.1 and offer practical advantages that improve alignment performance.
+Genetic algorithms maintain a population of candidate solutions, $\psi_{i}$, and iteratively improve them via selection, crossover, and mutation [eiben2015introduction].
+In Eq.˜2, the search distribution is an empirical distribution over the population of solutions, as shown in Fig.˜3.
+GAs offer practical advantages that improve alignment performance.
 First, the optimization procedure—selection, crossover, mutation—under correct parameterization can ensure the noise search is regularized to the high-density shell, provided the initial population resides within.
-In particular, uniform crossover (independent coordinate swaps) or permutation-based mutations do so—see § B.4 for details.
-Second, maintaining a population enables broad, parallel exploration early in optimization, while increasing selection pressure later drives convergence towards the promising regions (§ D.4).
-This capacity to explore multiple basins simultaneously makes GAs well-suited to multi-modal alignment objectives, e.g., human preferences.
-However, our experiments show that diversity collapse can cause premature convergence, hence marginal improvements over longer optimization timelines (see § 4.3 and § 4.4).
+In particular, uniform crossover (independent coordinate swaps) and permutation-based mutations do so.
+[^1] See our proof in Sec.
+B.5 for details.
+Second, maintaining a population enables broad, parallel exploration early in optimization, and selection pressure later (Sec.˜D.4).
+This makes GAs well-suited to multimodal alignment objectives, e.g.
+human preferences.
+However, we find that GAs can suffer from low sample diversity, leading to poor alignment over long optimization horizons (Secs.˜4.6 and 4.3).
 Natural Evolutionary Strategies (ES).
-Natural evolutionary strategies perform black-box optimization by adapting the parameterized search distribution $q_{\phi}$ over solutions [50].
-Unlike GAs, $q_{\phi}$ is a parameterized search distribution (e.g., multivariate normal).
-In the context of alignment, ES satisfy the criteria in § 3.1 and offer a distinct set of advantages over GAs.
-They maintain a search distribution, so their memory footprint is lower: only distribution parameters (e.g., mean, covariance, step size) need be stored.
+Natural evolutionary strategies perform black-box optimization by adapting the parameterized search distribution $q_{\phi}$ over solutions [wierstra2014natural].
+Unlike GAs, $q_{\phi}$ for ES is a parameterized search distribution (e.g.
+multivariate normal), depicted in Fig.˜3.
+ES offer different tradeoffs than GAs.
+They maintain a search distribution, so their memory footprint is lower: only distribution parameters (e.g.
+mean, covariance, step size) need be stored.
 This parameterized distribution also lets us draw an arbitrary number of candidate solutions, making ES ideal when many aligned samples are required.
-Moreover, although termed “gradient-free”, evolutionary strategies update their search distribution with a gradient approximation [36, 41].
-As a result, they move through the reward landscape with a coarse sense of direction lacked by GAs.
-This directional information translates into alignment improvements over long optimization horizons (see § 4.3).
-Choosing Between GA and ES for Alignment.
-Although GA and ES methods satisfy the criteria outlined in § 3.1, they exhibit markedly different properties.
-First, the search distribution for GA is an empirical distribution, whereas ES use multivariate normal distributions.
-Consequently, GA exhibit higher memory usage than ES because the entire population must reside in memory (see § 4.5).
-This has consequences for sampling as well: additional samples outside the population cannot be acquired for GA, whereas for ES, it merely requires sampling the parameterized distribution.
-Conversely, GA handle multi-modal rewards better than ES due to the initialization requirements of each EA (§ 3.4).
-However, we note that GAs often suffer from premature convergence whereas ESs do not (§ 4.4).
-3.4 Implementation and Computational Considerations
-This section discusses practical considerations for implementing inference-time alignment using evolutionary algorithms, focusing on initialization strategies and parallel evaluation.
+Moreover, although ES are “gradient free” in the sense of no backpropagation through network parameters, they update their search distributions with a gradient approximation [salimans2017evolution, such2017deep].
+Thus, ES move through reward landscapes with a sense of direction that GAs lack, resulting in improvements over long optimization horizons (Sec.˜4.6).
+Figure 3: We depict how GA and ES search the latent noise space (a Gaussian hypersphere [gaussian_annulus_lecture_2017]) over optimization steps $t$.
+GAs maintain an empirical population of solutions $q_{\phi}=\{\psi_{0}...\psi_{N}\}$ while ES maintains a distribution $q_{\phi}=\mathcal{N}(\mu_{i},\sigma_{i})$ from which we sample.
+3.3 Implementation Considerations
+This section discusses practical considerations for implementing evolutionary algorithms for alignment.
 Initialization and Sample Quality.
@@
 For GAs, initializing the population from the standard Gaussian distribution $\mathcal{N}(0,I)$ ensures that the initial samples reside within the high-density shell of the latent space.
-However, this initialization strategy does not translate to natural evolutionary strategies (ES), as subsequent updates to the search distribution can degrade sample quality.
+However, this strategy does not translate to ES, as subsequent updates to the search distribution can degrade sample quality.
 To mitigate this, we instead initialize $\mu$ of the search distribution to a fixed latent vector $z_{0}\sim\mathcal{N}(0,I)$ and use a small initial standard deviation.
-We elaborate further on consequences of ES initialization in Appendix C.
+Appendix˜C elaborates on the consequences of ES initialization.
 Parallel Evaluation and Population Scaling.
-Both GA and ES approaches evaluate a population of $N$ candidate solutions at each iteration.
-To improve runtime efficiency, we partition each population into batches of size $B$ and evaluate the batched solutions in parallel.
-Batching takes advantage of modern GPU parallelism and allows us to scale evaluations across larger populations without incurring excessive runtime costs.
-However, this comes at the cost of higher memory usage during model inference.
-In practice, the ability to choose the batch size for evaluation allows for a trade-off between memory usage and runtime.
-We explore this further in § 4.5.
+GA and ES approaches evaluate a set of $N$ candidate solutions at each iteration.
+To improve runtime efficiency, we partition $N$ solutions into batches of size $B$ and evaluate the batched solutions in parallel.
+This trades higher memory usage for lesser runtime costs due to inefficient serial processing.
+Detailed performance analysis is provided in Sec.˜4.5.
+Composability.
+Our evolutionary methods are compatible with other methods.
+Direct or transform noise search only modifies the initial noise $z_{T}$.
+Methods that do not target $z_{T}$ can be used with ours.
+We demonstrate this by composing our method with a fine-tuning approach (Sec.˜4.4).
 4 Evaluation
-We evaluate our noise optimization design decisions, and compare with other inference-time alignment work.
-§ 4.1 describes our experimental setup.
-§ 4.2 compares the ability of our method to maximize rewards with other work, notably Direct Noise Optimization [26] on the DrawBench and Open Image Preferences datasets.
-§ 4.3 investigates the relationship between optimization steps and alignment.
-§ 4.4 compares population statistics between ES and GA methods.
-§ 4.5 reports the memory consumption and runtime of our method across population size and batch size.
+We compare our EA alignment methods with other state-of-the-art inference-time alignment work.
+Sec.˜4.1 describes our experimental setup.
+Sec.˜4.2 compares the ability of our method to maximize rewards with other works.
+Sec.˜4.3 compares population statistics between ES and GA methods.
+Sec.˜4.4 investigates the effect of performing inference-time alignment on top of finetuning-based alignment.
+Sec.˜4.5 reports the memory consumption and runtime of our method across population size and batch size, including an equal-runtime evaluation.
+Sec.˜4.6 investigates the relationship between optimization steps and alignment.
 4.1 Experimental Setup
-We evaluate our alignment methods on a variety of diffusion models, benchmarks, and evolutionary algorithms. All experiments are conducted on NVIDIA A100 80GB GPUs.
-Evolutionary Algorithms We use a variety of state-of-the-art evolutionary algorithms: CoSyne, PGPE, and SNES, as provided in EvoTorch [44].
-Diffusion Models We implement EA-based search on three state-of-the-art diffusion models: Stable Diffusion-1.5 [34], Stable Diffusion 3 [9], and the latent consistency variant (LCM) of PixArt-$\alpha$ [3].
-Our choices reflect the diversity of diffusion model architectures and training methods.
-Datasets We follow [26], and conduct quantitative evaluations on DrawBench [35].
-DrawBench consists of 200 prompts from 11 categories that test the semantic properties of model, such as their ability to handle composition, cardinality, and rare prompts.
-We conduct our ablation studies and qualitative evaluations on a subset (60 prompts, 5 per category) of the Open Image Preferences dataset [1].
-Open Image Preferences consist of detailed and diverse prompts along with human preferences, which can be used to fine-tune diffusion models [6].
-For our purposes, both datasets provide complex prompts to evaluate the quality of inference-time methods.
-Reward Functions In accordance with other diffusion works [42, 23], we use ImageReward [53], CLIP [32], HPSv2 [51], and JPEG compressibility as reward functions and evaluation metrics.
+We evaluate three EAs on two prompt datasets and four reward functions. All experiments are conducted on one NVIDIA A100 80GB (PCIe) GPU.
+Algorithms.
+For our approach, we use three state-of-the-art EAs: CoSyNE, PGPE, and SNES (as provided in EvoTorch [toklu2023evotorch]).
+In terms of state-of-the-art methods: we use the gradient-based approach, DNO [tang2024inference_dno]; and three gradient-free approaches — SVDD [li_svdd_2024], DSearch-R [li2025dynamic], and Fk Steering (FKS) [singhal_fk_steering_2025].
+As baseline methods, we follow [ma2025inference_time_scaling] by using Best-of-N search (sample $N$ images, and keep the best reward sample) and Zero-Order search (sample $N$ initial latent noises, and sample using the one with the best reward).
+Optimization Horizons.
+We evaluate our methods and DNO using optimization horizons of length 15 (“short”) or 50 (“long”), similar to DNO [tang2024inference_dno] which evaluates over horizons of length 10, 50, and 100.
+We configure SVDD, DSearch-R, and FKS such that they have roughly equal runtime per prompt as our method, since they do not have a notion of optimization steps or population size in the same sense as our methods (cf.
+Sec.˜C.2).
+Diffusion Models.
+Here, we perform all evaluations on Stable Diffusion 1.5 [rombach2022high].
+The Appendix contains results for Stable Diffusion 3 [esser2024scaling] and the latent consistency variant (LCM) of PixArt-$\alpha$ [chen2023pixart].
+Prompt Datasets.
+We use two prompt datasets: DrawBench [drawbench2022] and Open Image Preferences [huggingface_image_preferences].
+Both datasets provide complex prompts to evaluate the quality of inference-time methods.
+DrawBench comprises 200 prompts from 11 categories that test the semantic properties of model, such as their ability to handle composition, cardinality, and rare prompts.
+Regarding Open Image Preferences, it features detailed prompts used to fine-tune diffusion models [data_is_better_together_open_image_preferences_v1_flux_dev_lora_2024].
+For additional ablation evaluations, we use a randomly sampled subset (60 prompts total, 5 per category) of the full dataset.
+Reward Functions and Measurement.
+Following prior diffusion works [tang2024inference_dno, li2025dynamic], we experiment with four rewards for alignment: ImageReward [xu2023imagereward], CLIP [clip2021], HPSv2 [hpsv2wu2023human], and JPEG size.
+JPEG size is a proxy for compressibility, while the other three metrics evaluate image aesthetic quality for inter-method comparison.
+In reporting reward statistics, we average over the dataset length, or number of prompts.
+For granular prompt-level reward analysis, see Appendix˜D.
+Table 1: DrawBench Evaluation.
+We evaluate various algorithms on four popular alignment objectives.
+Algorithms (CoSyNE, SNES, PGPE) are our evolutionary methods.
+CoSyNE achieves the highest rewards across all metrics, except one evaluation.
+We conduct search using each algorithm for a short-optimization horizon (15 steps) with a population of 16 solutions, except for DNO which has a population of 1.
+Best-of-N was conducted with N=240.
+N.B.
+$\uparrow$$\downarrow$ notation indicates a goal of maximization or minimization, respectively.
+Red indicates evidence of reward hacking.
+Algorithm ($\uparrow$) ImageReward ($\uparrow$) CLIP ($\uparrow$) HPSv2 ($\downarrow$) JPEG Size (kB)
+Best Sample Mean Best Sample Mean Best Sample Mean Best Sample Mean
+Mean $\pm$ Std Sample Mean $\pm$ Std Sample Mean $\pm$ Std Sample Mean $\pm$ Std Sample
+Best-of-N $1.41\pm 0.52$ $0.38$ $37.1\pm 3.7$ $32.9$ $0.301\pm 0.02$ $0.282$ $66.9\pm 17.8$ $105.4$
+Zero‑Order $1.46\pm 0.53$ $0.97$ $37.6\pm 3.9$ $34.5$ $0.304\pm 0.02$ $0.290$ $53.3\pm 18.4$ $62.8$
+DNO $0.71\pm 0.93$ - $26.2\pm 3.6$ - $0.286\pm 0.02$ - $58.1\pm 17.5$ -
+FKS $1.46\pm 0.51$ $1.25$ $28.7\pm 3.6$ $32.2$ $0.284\pm 0.02$ $0.279$ $62.6\pm 19.3$ $77.4$
+SVDD $1.04\pm 0.74$ $0.26$ $6.4\pm 3.9 33$.633.6 $0.292\pm 0.02$ $0.281$ $40.3\pm 16.9$ $54.1$
+DSearch-R $0.86\pm 0.89$ $0.72$ $36.2\pm 4.2$ $35.5$ $0.289\pm 0.02$ $0.286$ $37.8\pm 16.1$} 38.1{38.1}
+Ours: Direct Noise Search
+CoSyNE 1.61 $\pm$ 0.42 1.38 38.8 $\pm$ 3.8 36.5 0.310 $\pm$ 0.02 0.300 $39.5\pm 15.9$} 44.8{44.8}
+SNES $1.39\pm 0.52$ $0.78$ $38.1\pm 3.8$ $35.6$ $0.300\pm 0.02$ $0.286$ $71.3\pm 21.4$ $78.0$
+PGPE $1.26\pm 0.65$ $0.92$ $37.1\pm 3.6$ $34.4$ $0.299\pm 0.02$ $0.290$ $88.3\pm 23.7$ $97.0$
+Ours: Noise Transform Search
+CoSyNE $1.53\pm 0.44$} 1.23{1.23} $38.4\pm 3.8$} 36.5{36.5} $0.307\pm 0.02$} 0.299{0.299} $38.7\pm 16.3$ $42.3$
+SNES $1.34\pm 0.56$ $0.94$ $\pm 3.8 34.734.7$ $0.300\pm 0.02$ $0.290$ $57.8\pm 22.4$ $66.1$
+PGPE $1.28\pm 0.58$ $0.88$ $36.9\pm 3.8$ $34.8$ $0.300\pm 0.02$ $0.290$ 18.3 $\pm$ 12.4 20.2
+Table 2: Open Image Preferences Evaluation.
+We evaluate various methods on ImageReward and JPEG compression, configuring in the same manner as in Tab.˜1.
+Our method based on CoSyNE achieves the highest rewards on ImageReward, while achieving the third-lowest JPEG size.
+Algorithm ($\uparrow$) ImageReward ($\downarrow$) JPEG Size (kB)
+Best Sample Mean Best Sample Mean
+Mean $\pm$ Std Sample Mean $\pm$ Std Sample
+Best-of-N 1.49 $\pm$ 0.37 0.49 72.6 $\pm$ 23.5 116.2
+Zero–Order 1.53 $\pm$ 0.35 0.98 55.3 $\pm$ 21.3 66.7
+DNO 0.78 $\pm$ 0.76 - 61.2 $\pm$ 17.4 -
+FKS 1.54 $\pm$ 0.38 1.35 64.5 $\pm$ 19.4 79.6
+SVDD 1.06 $\pm$ 0.64 0.46 44.1 $\pm$ 18.2 59.2
+DSearch-R 1.01 $\pm$ 0.66 0.85 40.9 $\pm$ 17.7 41.3
+Ours: Direct Noise Search
+CoSyNE 1.71 $\pm$ 0.24 1.47 44.2 $\pm$ 14.6 49.7
+SNES 1.56 $\pm$ 0.37 1.22 83.3 $\pm$ 26.1 90.8
+PGPE 1.42 $\pm$ 0.40 0.94 98.1 $\pm$ 34.7 110.3
 4.2 Cross-Dataset Evaluation
+Tab.˜1 summarizes the results on DrawBench, and Tab.˜2 on Open Image Preferences. To assist interpretation, we also discuss reward hacking.
 DrawBench.
-Tab.
-1 depicts the efficacy of EAs at alignment.
-After only 15 optimization steps, EAs outperform both baseline black-box methods (Random, Zero-Order) and the gradient-based DNO on all four reward functions, for both the best sample and the median reward in the population.
-When searching over noise, CoSyNE delivers the strongest overall performance, achieving the best alignment performance on ImageReward, CLIP, HPSv2 scores and yields the largest JPEG-size reduction.
-The same pattern holds when we search over noise transformations, with CoSyNE again leading three of the four metrics; the sole exception is the JPEG-compression objective, where PGPE attains the smallest file sizes.
-Manual inspection, however, reveals that many of PGPE’s images omit key prompt content, suggesting reward hacking (see § D.5.1 for examples).
-Crucially, both EAs and the Zero-Order baseline improve median as well as best rewards, whereas Random search does not — indicating that directed search can steer us toward parts of the latent space where aligned samples reside.
-We elaborate on the consequences of searchable latent spaces in § 5.
-Table 1: DrawBench Evaluation on StableDiffusion-1.5 We show the best-sample reward and median reward of the best population achieved by each algorithm on DrawBench under four reward functions.
-Algorithms (Random, Zero-Order, DNO) are baselines, while (CoSyNE, SNES, PGPE) are evolutionary methods.
-We conduct search using each algorithm for 15 steps with a population of 16 solutions, except for DNO which has population of 1.
-NB: $\uparrow$/$\downarrow$ notation indicates a goal of reward maximization or minimization, respectively.
-Entries marked with a † are statistically insignificant ($p$ > 0.05), otherwise they are statistically significant.
-Algorithm ($\uparrow$) ImageReward ($\uparrow$) CLIP ($\uparrow$) HPSv2 ($\downarrow$) JPEG Size (kB)
-Best Median Best Median Best Median Best Median
-Random $1.41$ $0.38$ $37.1$ $32.9$ $0.301$ $0.282$ $66.9$ $105.4$
-Zero-Order $1.46$ $0.97$ $37.6$ $34.5$ $0.304$ $0.290$ $53.3$ $62.8$
-DNO $0.71$ $26.2$ $0.286$ $94.6$
-Ours: Noise
-CoSyNE $\mathbf{1.61}$ $\mathbf{1.38}$ $\mathbf{38.8}$ $\mathbf{36.5}$ $\mathbf{0.310}$ $\mathbf{0.300}$ 39.5 44.8
-SNES $1.39$ $0.78$ $38.1$ $35.6$ $0.300$ $0.286$ $71.3$ $78.0$
-PGPE $1.26$ $0.92$ $37.1$ $34.4$ $0.299$ $0.290$ $88.3$ $97.0$
-Ours: Noise Transformation (QR)
-CoSyNE $\mathbf{1.53}$ $\mathbf{1.23}$ $\mathbf{38.4}$ $\mathbf{36.5}$ $\mathbf{0.307}$ $\mathbf{0.299}$ $38.7$ $42.3$
-SNES $1.34$ $0.94$ $37.4$ $34.7$ $0.300$ $0.290$ $57.8$ $66.1$
-PGPE $1.28$ $0.88$ $36.9$ $34.8$ $0.300$ $0.290$ $\mathbf{18.3}$ $\mathbf{20.2}$
-Table 2: Open Images Preferences Cross-Model Evaluation The best-sample reward and the median reward of the best population achieved by each algorithm on Open Image Preferences across three models: StableDiffusion 1.5, StableDiffusion 3, and PixArt-$\alpha$.
-Algorithms are configured in the same manner as in  Tab.
-1.
-Higher rewards are better.
-Algorithm SD1.5 SD3 PixArt-$\alpha$ (LCM)
-CLIP ImgReward CLIP ImgReward CLIP ImgReward
-Best Med. Best Med. Best Med. Best Med. Best Med. Best Med.
-Random $39.1$ $34.9$ $1.49$ $0.59$ $39.0$ $36.0$ $1.72$ $1.38$ $38.8$ $35.9$ $1.67$ $1.25$
-Zero–Order $39.5$ $36.4$ $1.54$ $0.98$ $39.4$ $37.4$ $1.75$ $1.60$ $38.9$ $36.1$ $1.67$ $1.29$
-Ours: Noise
-CoSyNE $\mathbf{40.7}$ $\mathbf{38.3}$ $\mathbf{1.69}$ $\mathbf{1.47}$ $\mathbf{40.6}$ $\mathbf{39.0}$ $\mathbf{1.82}$ $\mathbf{1.75}$ $\mathbf{39.2}$ $\mathbf{36.8}$ $\mathbf{1.69}$ $\mathbf{1.45}$
-SNES $40.1$ $37.6$ $1.57$ $1.25$ $39.9$ $38.1$ $1.77$ $1.68$ $38.9$ $35.8$ $1.68$ $1.21$
-PGPE $39.0$ $36.1$ $1.37$ $0.88$ $38.4$ $36.4$ $1.67$ $1.50$ $38.7$ $35.8$ $1.65$ $1.23$
+Tab.˜1 depicts our results on DrawBench.
+Given a short optimization horizon (15 steps), EAs outperform both baseline black‑box methods (Best-of-N, Zero‑Order) and the gradient‑based DNO on all four reward functions, for both the best sample and the mean reward in the population.
+When searching over noise, CoSyNE delivers the strongest performance, achieving the best alignment performance on ImageReward, CLIP, and HPSv2 scores, while DSearch-R attains the lowest JPEG size.
+The same pattern holds when we search over noise transformations, with CoSyNE again leading three of the four metrics; the sole exception is the JPEG‑compression objective, where PGPE attains the smallest file sizes (however, see subsequent discussion of reward hacking).
 Open Image Preferences.
-Tab.
-2 analyzes three diffusion models and confirms that evolutionary algorithms largely remain competitive across models.
-For both SD1.5 and SD3, the EA family consistently surpasses the black-box baselines: CoSyNE achieves the highest best and median rewards on every metric, e.g., SD3-CLIP rises from 39.4/37.4 with Zero-Order to 40.6/39.0, while SD3-ImageReward improves from 1.75/1.60 to 1.82/1.75.
-SNES and PGPE follow the same trend, but with smaller margins.
-Generalization is weaker on PixArt-$\alpha$, where improvements vanish, suggesting that black-box latent search may not transfer to architectures trained with alternate objectives such as LCM — we elaborate in § 5.
-Overall, we show evolutionary search can align multiple diffusion models; however some latent spaces may be more difficult to search than others.
+Tab.˜2 depicts our results on Open Image Preferences.
+Similar to DrawBench (Tab.˜1), CoSyNE is the best performing method under equal (15) optimization steps.
+CoSyNE achieves the best alignment performance in terms of best and mean rewards on ImageReward and JPEG compression, and outperforming our other variants SNES and PGPE.
+It achieves the highest ImageReward scores across methods, and achieve lower but similar JPEG scores to SVDD and DSearch-R.
+Notably, our ImageReward scores are significantly higher than white-box methods DNO, FKS, SVDD, DSearch-R.
+Reward Hacking.
+To assess the incidence of reward hacking, we audited the best method for each alignment objective, and made note of obvious cases of any reward hacking across all of its images.
+On DrawBench, PGPE (ours), consistently produced images indicative of reward hacking on JPEG compression, with monochromatic results lacking prompt details.
+We observed minor or one-off instances of reward hacking on other methods, which we visualize in Sec.˜D.5.1.
+4.3 Population Statistics
+Figure 4: Reward Standard Deviation per Step (Diversity).
+CoSyNE (GA) has rapidly diminishing diversity, while SNES (ES) maintains diversity over optimization steps, as its search distribution evolves (Sec.˜3.2).
+This implies that GAs are suited for short-term optimization, and ES for long-term optimization.
+N.B.
+We use standard deviation as a proxy for solution diversity.
+Fig.˜4 plots the population’s reward standard deviation over 50 optimization steps for CoSyNE (a GA) and SNES (an ES), confirming the dynamics predicted in Sec.˜3.2.
+CoSyNE begins with high variance, reflecting its broad exploration capability, but diversity collapses under selection pressure (more ablations in Sec.˜D.4).
+In contrast, SNES sustains a moderate level of variance throughout optimization, preserving exploratory capacity even at later steps.
+This sustained diversity underpins the long-horizon gains of ES (Fig.˜6(a)).
+Taken together, these results suggest that GAs are ideal for rapid, short-horizon alignment; otherwise ES are preferable.
+4.4 Composition with Fine-tuning Alignment
+Table 3: Inference-Time Alignment and Finetuning-Based Alignment.
+We apply our inference-time alignment method on a fine-tuned model, Diffusion-DPO using ImageReward on the Open Image Preferences dataset.
+Our method improves rewards when used on top of this fine-tuning alignment method.
+cf: Tab.˜2.
+Model Inf. Time Best Sample Mean
+Algorithm Mean $\pm$ Std Sample
+SD1.5 Best-of-N 1.49 $\pm$ 0.37 0.49
+Diffusion-DPO Best-of-N 1.54 $\pm$ 0.34 0.66
+SD1.5 CoSyNE 1.71 $\pm$ 0.24 1.47
+SD1.5 SNES 1.56 $\pm$ 0.37 1.27
+Diffusion-DPO CoSyNE 1.74 $\pm$ 0.22 1.55
+Diffusion-DPO SNES 1.61 $\pm$ 0.32 1.31
+As noted in Sec.˜3.3, our method is compatible with fine-tuning approaches.
+Thus, we evaluate how our inference-time approach composes with a fine-tuning approach, Diffusion-DPO [wallace2024diffusion] on ImageReward.
+Tab.˜3 depicts reward statistics of our approach applied to Diffusion-DPO finetuned Stable Diffusion 1.5, and the standard Stable Diffusion 1.5.
+It can be seen that CoSyNE achieves higher best and mean rewards in composition with Diffusion-DPO, indicating our method can be combined with finetuning based methods to further increase alignment.
+(a) Memory usage vs. batch size (PP=$B$) on A100 (80GB VRAM).
+(b) Running time per step vs. batch size. (PP=16)
+Figure 5: Computational characteristics of inference‑time alignment methods on Stable Diffusion-1.5.
+In terms of memory usage our evolutionary methods scale with batch size, whereas DNO does not and exhausts memory for batch sizes $\geq 16$.
+Our methods exhibit near-constant or decreasing per‑step latency as batch size increases, indicating that they can be efficiently batched.
+For DNO, $P=B$.
+(a) Best-sample reward per step on ImageReward.
+(b) Mean reward per step on ImageReward.
+Figure 6: Reward per-step measured on Open Image Preferences.
+Evolutionary methods (CoSyNE, SNES, PGPE) outperform Zero‐Order and Best-of-N baselines.
+Best-of-N was run with N=800.
+Algorithms were run with population size PP=16, except DNO with PP=1.
+4.5 Computational Costs
+We discuss computational costs of surveyed methods in terms of runtime per prompt and memory consumption.
+[…]
+Table 4: Runtime Evaluation with ImageReward.
+We use runtime as a proxy for compute, evaluating each method on Open Image Preferences with ImageReward over 50 optimization steps.
+CoSyNE achieves the highest reward for similar, and in some cases significantly more, runtime across all methods and baselines.
+Methods are configured to closely match our runtime, except Best-of-N (N=2500) to exemplify its sample inefficiency.
+Algorithm Best Sample Runtime
+Mean Per Prompt (s)
+Best-of-N 1.37 2104
+Zero-order 1.66 1343
+DNO 1.26 1270
+FKS 1.56 1489
+SVDD 1.72 1180
+DSearch-R 1.15 1009
+CoSyNE (PP=16) 1.78 1119
+CoSyNE (PP=10) 1.57 814
+SNES (PP=16) 1.74 1485
+PGPE (PP=24) 1.68 1126
+Batching Memory Comparison.
+We compare the memory usage and runtime per step for evolutionary methods (CoSyNE, PGPE, SNES) in Sec.˜3 with Best-of-N, Zero-Order search, and DNO.
+We measure runtime per optimization step and peak GPU memory usage with respect to batch size ($B$) and population size ($P$), as stated in Sec.˜3.3.
+Fig.˜5(a) shows that as batch size $B$ increases, evolutionary methods consume significantly less GPU memory than DNO (gradient-based).
+DNO consumes $2.2\times$ to $4.1\times$ more GPU memory than EAs, and runs out-of-memory at batch sizes $\geq 16$, whereas evolutionary methods do not.
+Compared with other gradient-free approaches Best-of-N and Zero-Order search, EAs consume $1\times$ to $1.5\times$ more GPU memory, but achieve better alignment per earlier results (Sec.˜4.2).
+Fig.˜5(b) shows the running time per step for each method across batch size $B$.
+All EAs exhibit a decrease in runtime per step as $B$ grows: at small $B$, the population of $P$ candidates must be evaluated sequentially, thus runtime is high.
+As $B$ increases, evaluations are batched and parallelized, reducing the time per step at the expense of additional memory usage as shown in Fig.˜5(a).
+In contrast, DNO’s per-step runtime increases with $B$ and eventually exceeds all methods.
 4.6 Optimization Steps and Alignment
-We evaluate how the number of inference-time optimization steps affects alignment.
-Fig.
-3 depicts the best-sample and median reward per optimization step on Open Image Preferences benchmark, using ImageReward alignment objective.
-All methods improve their best-sample and median rewards as the number of steps increases, but they differ in step-efficiency.
-The population-based EA Cosyne achieves the highest step-efficiency, yielding the highest best-sample and median reward at every step count.
-Whereas, the distribution-based EAs, SNES and PGPE, require more steps to surpass random and zero-order search.
-SNES needs more 15 than steps to outperform zero-order in best-sample performance.
-Meanwhile, PGPE needs more than 10 steps to exceed random search in best-sample performance.
-When considering median reward, all EAs exceed baselines given enough optimization steps.
-(a) Best-sample reward per step on ImageReward.
-(b) Median reward per step on ImageReward.
-Figure 3: Reward per step measured on Open Image Preferences.
-Evolutionary methods (Cosyne, SNES, PGPE) outperform Zero-Order and Random baselines.
-Each algorithm optimizes the reward for 50 steps with a population of 16, except for DNO which has a population of 1.
-4.4 Population Statistics
-Fig.
-4 plots the population’s reward standard deviation over 50 optimization steps for CoSyNE (a GA) and SNES (an ES), confirming the dynamics predicted in § 3.3.
-CoSyNE begins with high variance, reflecting broad exploration, but diversity collapses under selection pressure (§ D.4 gives ablations on selection pressures).
-Figure 4: Reward standard deviation per step, measured on Open Image Preferences.
-CoSyne (GA) has high diversity early, but this rapidly diminishes.
-SNES (ES) has a more uniform pattern, as its search distribution adapts per step (§ 3.3).
-NB: Reward standard deviation proxies for solution diversity.
-In contrast, SNES sustains a moderate level of variance throughout optimization, preserving exploratory capacity even at later steps.
-This sustained diversity underpins the long-horizon gains of ES (Fig.
-3(a)).
-Taken together, these results suggest that GAs are ideal for rapid, short-horizon alignment, while ES are preferable when more time is available.
-4.5 Computational Costs
-We compare the memory usage and runtime per step for evolutionary methods in § 3 with Random, Zero-Order Search, and DNO.
-We measure the runtime per optimization step and the peak GPU memory usage with respect to batch size ($B$) and population size ($P$), as stated in § 3.4.
-Fig.
-5(a) shows that as batch size $B$ increases, evolutionary methods consume significantly less GPU memory than DNO.
-The gradient-based DNO consumes $\times2$ to 4.1×4.1$\times$4.1 × more GPU memory than evolutionary methods.
-Moreover, DNO runs out of memory (OOM) at batch sizes $\geq 16$, whereas evolutionary methods do not.
-Compared to other gradient-free approaches (Random and Zero-Order search), evolutionary methods consume only 1×1$\times$1 × to 1.5×1.5$\times$1.5 × more GPU memory.
-However, in terms of alignment performance (Tab.
-1), evolutionary methods outperform both Random and Zero-Order search with small additional memory costs.
-Fig.
-5(b) shows the running time per step for each method as we vary the batch size $B$.
-All evolutionary algorithms (CoSyne, PGPE, SNES) exhibit a decrease in runtime per-step as $B$ grows: at small $B$, the population of $P$ candidates must be evaluated sequentially, thus runtime is high.
-As $B$ increases, evaluations are batched and parallelized, reducing the time per step at the expense of additional memory usage as shown in Fig.
-5(a).
-In contrast, DNO’s per-step runtime increases with B and eventually exceeds all methods at large batch sizes.
-(a) Memory usage vs. batch size ($P$= $B$).
-(b) Running time per step vs. batch size. $(P=16)$
-Figure 5: Computational characteristics of inference-time alignment methods on Stable Diffusion-1.5.
-In terms of memory usage evolutionary methods (CoSyne, PGPE, SNES) scale with batch size, whereas DNO does not and exhausts memory for batch sizes $\geq 16$.
-All methods aside from DNO exhibit near-constant or decreasing per-step latency as batch size increases.
-For DNO: $(P=B)$
-5 Discussion and Limitations
-Searching Latent Spaces.
-By virtue of our results, we show that optimizing noise with EAs is often sufficient to perform alignment (see Tab.
-1), though its effectiveness varies by model (Tab.
-2).
-This raises two questions: (1) Are some latent spaces easier to search than others?
-(2) How can one design, train, or modify diffusion models to make their latent spaces easier to search?
-Our reward improvements on PixArt-$\alpha$ were diminished compared to other diffusion models (Tab.
-2), suggesting some property of the latent space and/or diffusion model that affects our search efficacy.
-Limitations.
-Our approach is ill-suited for long optimization horizons.
-With enough time, white-box methods that exploit gradient information will achieve better alignment (see § D.2).
-Black-box alignment fares better on short time-horizon applications, e.g., resource-constrained tasks such as mobile image generation.
-It is less suitable in contexts with high resource budgets, e.g., drug discovery.
-Even over short optimization horizons, our EA techniques shares the limit of other gradient-free and gradient-based methods (§ 2.1): some engineering time is incurred in tuning, and regularization is required to maintain sample quality (§ 3).
-We did not evaluate our method across modalities other than images, but we conjecture that our method could perform well across modalities.
+We evaluate how the number of inference‑time optimization steps affects alignment.
+Fig.˜6 depicts the rewards per optimization step on Open Image Preferences benchmark, using ImageReward as our alignment objective.
+All methods improve their rewards as the number of steps increases, but they differ in step‑efficiency, or reward increase per step.
+CoSyNE (a GA) achieves the highest step-efficiency, yielding the highest best-sample and mean reward at every step count.
+Meanwhile, the EA-based methods exceed baseline mean rewards with enough steps.
+PGPE needs >>10 steps to exceed Best-of-N in best-sample performance and SNES needs >>15 steps to beat Zero-Order search.
+5 Limitations
+Black-box methods and EAs are likely outperformed by white-box methods in terms of pure reward maximization over long optimization horizons.
+Although CoSyNE performed better on short optimization horizons (Tabs.˜1 and 2) than white-box approaches, its advantage diminishes in long term optimization (Secs.˜D.2 and 8).
+This suggests our EA-based methods are less suitable in contexts where reward maximization is the only concern, but may be preferable when seeking high sample-efficiency.
 6 Conclusion
 We illustrate how inference-time alignment for diffusion models can be performed with evolutionary algorithms (EA).
-Specifically, we search for the initial noise vector used in the reverse denoising process in order to alter the sampled image and maximize some reward.
-Our results illustrate that EAs can achieve higher rewards than existing inference-time alignment methods across a set of various diffusion models and datasets.
-In terms of compute resource requirements, our EA-based method uses 55%-76% lower GPU memory and runs 72%-80% faster than gradient-based alignment methods.
-Our findings highlight the effectiveness of EAs for inference-time alignment via noise optimization.
+Specifically, we search for the initial noise vector used in the reverse denoising process to maximize alignment objectives.
+Our results illustrate that EAs can achieve higher rewards than existing inference-time alignment methods, particularly in short optimization horizons.
+Although EAs are sample-efficient, white-box methods likely have an advantage in long optimization horizons.
+A promising direction for future work may include designing or specializing EAs to handle longer horizons.
 Appendix A Overview of Appendices and Supplemental Material
-  • Appendix A: This summary.
-  • Appendix B: Extended commentary on related work § 2.1 and § 2.2, with insights and details on our implementation of evolutionary algorithms for alignment.
-  • Appendix C: Summary of insights we encountered when deciding how to initialize genetic algorithms and evolutionary strategies, related to § 3.3.
-  • Appendix D: Extended quantitative and qualitative results across various models, reward functions, datasets, and algorithms used in our work, complementary to § 4.2§ 4.4.
+  • Appendix˜A: This summary.
+  • Appendix˜B: Extended commentary on related work Sec.˜2.1 and Sec.˜2.2, with insights and details on our implementation of evolutionary algorithms for alignment.
+  • Appendix˜C: Summary of insights we encountered when deciding how to initialize genetic algorithms and evolutionary strategies, related to Sec.˜3.2.
+We include configuration details of compared methods such as SVDD, DSearch, and FkD for fair comparison.
+  • Appendix˜D: Extended quantitative and qualitative results across various models, reward functions, datasets, and algorithms used in our work, complementary to Sec.˜4.2Sec.˜4.3.
 Appendix B Extended Commentary on Related Work and Implementations
 Here, we provide additional detail for concepts in our work that was not able to be fit into the main text.
+We include a small table of qualitative differencesies between alignment methods ( Tab.˜5), which complements  Sec.˜2.1.
 B.1 Extended Diffusion Model Alignment Related Work
-Diffusion models use a reverse diffusion process to convert some latent noise distribution into a data distribution, such as images [40, 16].
+Diffusion models use a reverse diffusion process to convert some latent noise distribution into a data distribution, such as images [sohl-dickstein_deep_2015, ho2022video_diffuion].
 The reverse diffusion process iteratively denoises the initial latent noise $z_{T}$ over some number of steps ($t=T\xrightarrow{}t=0$) to yield a sample, $z_{0}$.
@@
 Diffusion model alignment methods adjust diffusion models such that the resulting samples better meet an objective beyond the model’s original maximum likelihood criterion.
-These objectives commonly focus on producing images that reflect human aesthetic preferences [17, 51, 52, 53], or other metrics such as compressibility [2].
-Table 3: Comparison of inference-time diffusion alignment methods.
-Method Arbitrary Rewards Gradient Free Black-box
-DOODL [47]
-DNO [42]
-DSearch [23]
-Search-Paths [26]
-Zero-Order [26]
+These objectives commonly focus on producing images that reflect human aesthetic preferences [kirstain2023pick, hpsv2wu2023human, hpswu2023human, xu2023imagereward], or other metrics such as compressibility [black2023training].
+Table 5: Qualities of diffusion alignment methods.
+Method Arbitrary Gradient Black
+Rewards? Free? Box?
+DNO [tang2024inference_dno]
+Diffusion-DPO [wallace2024diffusion]
+DSearch-R [li2025dynamic]
+SVDD [li2025dynamic]
+FKS [singhal_fk_steering_2025]
+Zero-Order [ma2025inference_time_scaling]
+Best-of-N
 Ours
@@
 Fine-tuning-based alignment methods involve adjusting the diffusion model’s parameters so that generated samples better match alignment objectives.
-Generally, these method rely on curated datasets [37] or reward models [53, 52, 51] to fine-tune diffusion models [21, 5, 31, 4], and can involve supervised fine-tuning, or reinforcement learning based methods [48, 2].
+Generally, these method rely on curated datasets [laionAestheticDataset] or reward models [xu2023imagereward, hpswu2023human, hpsv2wu2023human] to fine-tune diffusion models [lee2023aligning, dai2023emu, prabhudesai2023aligning, clark2023directly], and can involve supervised fine-tuning, or reinforcement learning based methods [wallace2024diffusion, black2023training].
 Fine-tuning based methods, although powerful, require retraining and thus all the associated costs — our work does not require training.
@@
 Rather than altering model parameters, these techniques adjust sampling or conditioning to ensure samples meet alignment objectives.
-Formally, they seek the control variable $\psi$ that maximizes the expected reward $R(x)$ of samples $x$ from a pretrained diffusion model $p_{\theta}$, as shown in Eq.
-1.
+Formally, they seek the control variable $\psi$ that maximizes the expected reward $R(x)$ of samples $x$ from a pretrained diffusion model $p_{\theta}$, as shown in Eq.˜1.
 $\psi^{\star}=\underset{\psi\in\Psi}{\arg\max}\;\mathbb{E}_{x\sim p_{\theta}(x\mid\psi)}\bigl[R(x)\bigr].$ (3)
-Examples of control variables, $\psi$, include optimizing conditional input prompts [15, 12, 27, 10], manipulating cross-attention layers [11], and tuning latent noise vectors (noise optimization) to guide the diffusion trajectory [47, 42, 23, 45, 55, 26].
+Examples of control variables, $\psi$, include optimizing conditional input prompts [hao2023optimizing, gal2022image, manas2024improving, fei2023gradient], manipulating cross-attention layers [feng2022training], and tuning latent noise vectors (noise optimization) to guide the diffusion trajectory [wallace2023end, tang2024inference_dno, li2025dynamic, uehara2025inference, zhou2024golden, ma2025inference_time_scaling].
 In this work, we focus exclusively on noise optimization ($\psi=z$), which is generalizable across diffusion models.
@@
 Gradient-based methods refine the noise iteratively by leveraging the gradient with respect to the reward.
-Direct Optimization of Diffusion Latents, DOODL [47], and Direct Noise Optimization, DNO [42], are exemplary of these approaches.
+Direct Optimization of Diffusion Latents, DOODL [wallace2023end], and Direct Noise Optimization, DNO [tang2024inference_dno], are exemplary of these approaches.
 Both need to contend with the computational costs of backpropagation and maintaining sample quality by keeping the optimized noise on the Gaussian shell.
@@
 Likewise, DNO computes the gradient through the reward function to optimize the noise, but can optimize non-differentiable rewards by using zero-th order optimization algorithms.
-The drawbacks of gradient-based methods are: they are not black-box; they are computationally expensive, especially when aligning multiple samples; and often require long optimization budgets (e.g., DNO requires $>100$ optimization steps, whereas our method achieves better results in $\sim$50 steps).
+The drawbacks of gradient-based methods are: they are not black-box; they are computationally expensive, especially when aligning multiple samples; and often require long optimization budgets (e.g.
+DNO requires >100>100 optimization steps, whereas our method achieves better results in $\sim$50 steps).
 Gradient-free methods, on the other hand, explore the space of noise vectors or trajectories using only search or sampling methods.
-Ma et al.
-[26], employ three different strategies — random search, zero-order search (similar to hill climbing), and search-over-paths — to find samples that maximize their reward functions.
-Uehara et al.
-[45] provide a comprehensive overview of various inference-time algorithms, covering sequential Monte Carlo (SMC)-based guidance, value-based importance sampling, tree search, and classifier guidance.
+Ma et al[ma2025inference_time_scaling], employ three different strategies — random search, zero-order search (similar to hill climbing), and search-over-paths — to find samples that maximize their reward functions.
+Uehara et al[uehara2025inference] provide a comprehensive overview of various inference-time algorithms, covering sequential Monte Carlo (SMC)-based guidance, value-based importance sampling, tree search, and classifier guidance.
 Li et al.
-propose DSearch [23], a dynamic beam search algorithm in order to search for noise in order to maximize some reward function.
+propose DSearch [li2025dynamic], a dynamic beam search algorithm in order to search for noise in order to maximize some reward function.
 These techniques vary in their trade-offs, for example DSearch and the techniques outlined by Uehara et al.
 are not black-box and may have large computational costs.
-Meanwhile, although random and zero-order search are black-box and computationally efficient we show evolutionary methods outperform them (§ 4.2).
+Meanwhile, although random and zero-order search are black-box and computationally efficient we show evolutionary methods outperform them (Sec.˜4.2).
 In this work, we investigate a different class of gradient-free optimization methods: evolutionary algorithms.
@@
 B.2 Extended Alignment Objective Commentary
-Here, we make additional comments upon Eq. 2 as presented in § 3. Recall that Eq. 2:
+Here, we make additional comments upon Eq.˜2 as presented in Sec.˜3. Recall that Eq.˜2:
 $\phi^{\star}=\arg\max_{\phi}\;\underbrace{\mathbb{E}_{\psi\sim q_{\phi}(\psi)}}_{\begin{subarray}{c}\text{search}\\ \text{distribution}\end{subarray}}\Bigl[\mathbb{E}_{x\sim p_{\theta}(x\mid\psi)}\bigl[R(x)\bigr]\Bigr]\,$,
 cast inference-time alignment as search problem where we find the search distribution $q_{\phi}(\psi)$, that maximizes the expected reward.
-This objective is a generalization of Eq.
-1 to natural evolutionary strategies [50].
+This objective is a generalization of Eq.˜1 to natural evolutionary strategies [wierstra2014natural].
 Namely, we rather than searching for a single $\psi$ that maximizes the expected reward, we assume that $\psi$ is sampled from a parameterized search distribution, $q_{\phi}$, and we maximize the expected reward under this distribution.
@@
 Genetic Algorithms.
-Genetic algorithms maintain a population of candidate solutions, $\psi_{i}$, and iteratively improve them via selection, crossover, and mutation [8].
-In the context of Eq.
-2, the population is viewed as an empirical distribution.
+Genetic algorithms maintain a population of candidate solutions, $\psi_{i}$, and iteratively improve them via selection, crossover, and mutation [eiben2015introduction].
+In the context of Eq.˜2, the population is viewed as an empirical distribution.
 Thus the objective is to maximize the expected reward of a population of solutions.
@@
   1. 1.
-    Evaluation: For each solution $\psi_{i}$ in the population, the fitness (reward) is evaluated — i.e., $R(x_{i})$ where $x_{i}\sim p_{\theta}(x|\psi_{i})$.
+    Evaluation: For each solution $\psi_{i}$ in the population, the fitness (reward) is evaluated — i.e. $R(x_{i})$ where $x_{i}\sim p_{\theta}(x|\psi_{i})$.
   2. 2.
@@
 For example, when using uniform crossover (swapping each coordinate independently), each child’s coordinate is drawn from one of two i.i.d.
-$\mathcal{N}(0,1)$ parents thus remaining with high-density shell (§ B.4).
+$\mathcal{N}(0,1)$ parents thus remaining with high-density shell (Sec.˜B.5).
   4. 4.
@@
 Natural Evolutionary Strategies.
-Natural evolutionary strategies perform black-box optimization by adapting the parameterized search distribution $q_{\phi}$ over solutions [50].
-Unlike GAs, $q_{\phi}$ is a parameterized search distribution (e.g., multivariate normal) and at each iteration its parameters are updated as follows:
+Natural evolutionary strategies perform black-box optimization by adapting the parameterized search distribution $q_{\phi}$ over solutions [wierstra2014natural].
+Unlike GAs, $q_{\phi}$ is a parameterized search distribution (e.g.
+multivariate normal) and at each iteration its parameters are updated as follows:
   1. 1.
@@
 Thus, we update the parameters of the search distribution $q_{\phi}$ so that future samples are better aligned.
-B.4 Proof of Gaussian Marginal Preservation under Uniform Crossover
+B.4 Noise Transform Search Implementation
+Earlier we introduced the algorithm for direct noise search ( Algorithm˜1). We noted that the algorithm for noise noise transform search is slightly different: we include it here for clarity.
+Algorithm 2 Alignment via Noise Transformation Search
+1:Pretrained diffusion model $p_{\theta}$, reward $R(\cdot)$, noise $z_{T}\sim\mathcal{N}(0,I)$, iterations $T$, population size $M$
+2:Search variable: $\psi=A$, with transformed noise $z_{T}^{\prime}=Q(A)\,z_{T}$
+3: where $Q(A)$ denotes the orthonormal component of AA obtained via QR decomposition
+4:Search distribution $q_{\phi}(\psi)$:   GA: population $\{A_{i}\}_{i=1}^{M}$ or (ES) parameterized AA
+5:for $t=1,\dots,T$ do
+6:  Sample transformation parameters $\{A_{i}\sim q_{\phi}\}_{i=1}^{M}$
+7:  Compute transformed noises $\{z_{i}^{\prime}=Q(A_{i})\,z_{T}\}_{i=1}^{M}$
+8:  Generate samples $\{x_{i}\}_{i=1}^{M}$ with $x_{i}=f_{\theta}(z_{i}^{\prime})$
+9:  Compute rewards $\{r_{i}\}_{i=1}^{M}$ with $r_{i}=R(x_{i})$
+10:  $\phi\leftarrow\text{EAUpdate}(\phi,\{A_{i}\}_{i=1}^{M},\{r_{i}\}_{i=1}^{M})$
+11:end for
+12:return final $q_{\phi}$ or best transform
+NB: EAUpdate depends on the specific evolutionary algorithm used. A bias term $b$ may be included, but we set it to 0.
+B.5 Proof of Gaussian Marginal Preservation under Uniform Crossover
 Purpose
-In this section we show that uniform crossover maintains the solutions within the high-density Gaussian shell, which was alluded to in  § 3.3.
-This is relevant specifically to the EvoTorch implementation of the CoSyne algorithm, which uses a uniform distribution to decide crossover behavior.
-Claim
-Let $X=(X_{1},\dots,X_{d})$ and $Y=(Y_{1},\dots,Y_{d})$ be independent draws from $\mathcal{N}(0,I_{d})$. To perform coordinate-wise uniform crossover, for each $i$ we independently sample:
+In this section we show that uniform crossover maintains the solutions within the high-density Gaussian shell, which was alluded to in  Sec.˜3.2.
+This is relevant to the EvoTorch implementation of the CoSyne algorithm, which uses a uniform distribution to decide crossover behavior (parameterized by a probability $p$ to perform crossover).
+Background.
+Samples from a high-dimensional Gaussian lie on a hypersphere, this phenomenon is known as the Gaussian Annulus Theorem.
+Diffusion models that use Gaussian noise are (implicitly) trained under this condition [gaussian_annulus_lecture_2017].
+Thus, to ensure valid generated images, we must enforce this condition (draw samples near the surface of the sphere or shell) when we perturb noise.
+Uniform crossover is one such pertrubation, which we now investigate.
+Lemma.
+Let $X=(X_{1},\dots,X_{d})$ and $Y=(Y_{1},\dots,Y_{d})$ be independent draws from $\mathcal{N}(0,I_{d})$. To perform coordinate‐wise uniform crossover, for each $i$ we independently sample:
 $B_{i}\sim\mathrm{Bernoulli}\bigl(p\bigr),\quad Z_{i}=\begin{cases}X_{i},&B_{i}=1,\\ Y_{i},&B_{i}=0.\end{cases}$
 Then each coordinate $Z_{i}\sim\mathcal{N}(0,1)$ and thus $Z\sim\mathcal{N}(0,I_{d})$.
-Proof.
 Fix a coordinate $i$. For any real $z$, by the law of total probability conditioning on $B_{i}$:
-$P(Z_{i}\leq z)$ $=P(Z_{i}\leq z\mid B_{i}=1)\,P(B_{i}=1)+P(Z_{i}\leq z\mid B_{i}=0)\,P(B_{i}=0)$
-$=P(B_{i}=1)\,P(X_{i}\leq z)+P(B_{i}=0)\,P(Y_{i}\leq z)$
+$P(Z_{i}\leq z)$ $=P(Z_{i}\leq z\mid B_{i}=1)\,P(B_{i}=1)$
+$+P(Z_{i}\leq z\mid B_{i}=0)\,P(B_{i}=0)$
+$=P(B_{i}=1)\,P(X_{i}\leq z)$
+$+P(B_{i}=0)\,P(Y_{i}\leq z)$
 $=p\,\Phi(z)+(1-p)\,\Phi(z)$
@@
 $P(Z_{1}\leq z_{1},\dots,Z_{d}\leq z_{d})=\prod_{i=1}^{d}P(Z_{i}\leq z_{i})=\prod_{i=1}^{d}\Phi(z_{i}),$
-which is the CDF of $\mathcal{N}(0,I_{d})$. Equivalently, $Z\sim\mathcal{N}(0,I_{d})$. ∎
-Implication
-As noted in § 3.3, uniform crossover will ensure that solutions are with the high-density shell of the gaussian thus maintaining sample quality.
-Appendix C Details on Parameterization of Evolutionary Algorithms
-Here, we review more details on how we initialized our search parameters. We also comment on consequences and insights related to choices of initialization.
-GA Mutation Rate and Stability.
+which is the CDF of $\mathcal{N}(0,I_{d})$. Equivalently, $Z\sim\mathcal{N}(0,I_{d})$.
+Implication.
+As noted in Sec.˜3.2, uniform crossover will ensure that solutions are with the high-density shell of the gaussian thus maintaining sample quality.
+Appendix C Details on Parameterization of Evolutionary Algorithms and Compared Methods
+Here, we review more details on how we initialized our EA search parameters, and how other algorithms were configured for fair comparison.
+C.1 EA Parameterization
+Genetic Algorithm Mutation Rate and Stability.
 When using GAs, the mutation operator introduces Gaussian noise to promote exploration.
@@
 Here, we briefly describe pitfalls with improper ES initialization, and how we addressed them in our work.
-Fig.
-6 visually depicts this concept.
+Fig.˜7 visually depicts this concept.
 We noted that ES are particularly sensitive to initialization — in the context of this work, this "initialization" is the initial search distribution $q_{\phi}(\psi)$.
@@
 As a result, sample quality is degraded significantly.
-This is visually depicted in the top row of Fig.
-6.
+This is visually depicted in the top row of Fig.˜7.
 Figure 7: We depict a shortcoming with naive ES initialization, and illustrate our solution.
@@
 This modification was effective at improving sample quality, because our samples were now more closely drawn from the surface of the Gaussian shell.
-Appendix D Extended Evalaution Results
-Here, we introduce additional quantiative and qualitative results.
-D.1 Significance Testing
-We verify whether our results are statistically significant compared to unguided, random sampling of noise by performing Wilcoxon’s signed rank test; which assumes that distribution of pair differences is symmetric.
-We compare the resulting $p$ value from Wilcoxon’s against the (standard) significance level $\alpha=0.05$.
-Entries in  Tab.
-2 and  Tab.
-1 are annotated with a dagger if they are statistically insignificant.
-Across both tables, only 8 out of 84 of the entries related to our method (CoSyne, SNES, PGPE) were insignificant.
-D.2 Effect of longer optimization timelines for gradient-based methods
-Figures 7, 8 and 9 depict the effect of longer optimization timelines on the best-sample, mean, and median rewards for the gradient-based method DNO, compared to gradient-free methods.
+C.2 Configuration of Compared Methods
+Hyperparameter Description and Nuance.  We list the primary hyperparameters of each method in Tab.˜6; these define the configurations reported in Tab.˜7.
+Although some notions—such as “batch size” BB—appear across multiple methods, they do not necessarily correspond to the same underlying mechanism.
+In our method, $B$ either equals the population size ($B=P$) when evaluations are fully parallelized, or is smaller than the population ($B<P$) when we process individuals in batches (Sec.˜3.3).
+Other methods do not necessarily treat the batch dimension in the same manner.
+Instead, their batch of particles or samples is more similar to a population of samples to generate rather than the number of samples they evaluate per-step.
+For example, DSearch-R interprets $B$ as a number of samples to output — but it resamples and replaces samples within the batch throughout the denoising process.
+Choosing Configurations.
+Tab.˜7 shows configurations used in Tabs.˜1 and 2 .
+It was not straightforward to map our hyperparameters to those of other methods, given the differences noted above.
+To ensure a fair comparison, we therefore selected configurations that maximized alignment capability while matching the per-prompt running time of our strongest method, CoSyNE, on each reward function.
+Table 6: Hyperparameters of Surveyed Methods.
+This is a companion table for Tab.˜7.
+We refer the reader to the papers (and respective codebases) for a more detailed explanation of each.
+Batch size $B$ is interpreted differently by each method, per Sec.˜C.2.
+   Algorithm(s)    Parameter    Description
+   CoSyNE, SNES, PGPE    $P$    Population size
+   $B$    Batch size, parallel evaluations
+   $O$    Optimization Steps
+   FKS[singhal_fk_steering_2025]    PC    Particle Count, similar to population size $P$
+   $t_{\text{start}}$    Resampling denoising step start
+   $t_{\text{end}}$    Resampling denoising step end
+   $t_{\text{freq}}$    Resampling frequency
+   DSearch-R[li2025dynamic]    $B$    Similar to population size $P$
+   $n_{\text{duplicates}}$    Beam width
+   $w$    Tree expansion factor
+   SVDD[li_svdd_2024]    $B$    Similar to population size $P$
+   $n_{\text{duplicates}}$    Beam width
+Table 7: Tabs.˜1 and 2 Hyperparameter Configurations.
+These configurations closely match the running time of our best method, CoSyNE for each alignment objective.
+HPSv2 is slower to evaluate than all other alignment objectives, partially due to its large CLIP backbone.
+We note that batch size $B$ is interpreted differently by each method, per Sec.˜C.2.
+Algorithm(s) Objective Configuration
+FKS[singhal_fk_steering_2025] ImageReward, CLIP, JPEG $\text{PC}=128$, $t_{\text{start}}=5$, $t_{\text{end}}=45$, $t_{\text{freq}}=1$
+SVDD[li_svdd_2024] ImageReward, CLIP, JPEG $B=6$, $n_{\text{duplicates}}=20$
+DSearch-R[li2025dynamic] ImageReward, CLIP, JPEG $B=8$, $n_{\text{duplicates}}=5$, $w=5$, $r=0.125$
+FKS HPSv2 $\text{PC}=16$, $t_{\text{start}}=5$, $t_{\text{end}}=45$, $t_{\text{freq}}=1$
+SVDD HPSv2 $B=6$, $n_{\text{duplicates}}=10$
+DSearch-R HPSv2 $B=8$, $n_{\text{duplicates}}=4$, $w=4$, $r=0.125$
+CoSyNE, SNES, PGPE ImageReward, CLIP, HPSv2, JPEG $B=16$, $P=16$
+Appendix D Extended Evaluation Results
+D.1 Cross-Model Evaluation on Open Image Preferences
+Table 8: Open Images Preferences Cross-Model Evaluation The best‐sample reward and the median reward of the best population achieved by each algorithm on Open Image Preferences across three models: StableDiffusion 1.5, StableDiffusion 3, and PixArt-$\alpha$.
+Algorithms are configured in the same manner as in  Tab.˜1.
+Higher rewards are better.
+Algorithm SD1.5 SD3 PixArt-$\alpha$ (LCM)
+CLIP ImgReward CLIP ImgReward CLIP ImgReward
+Best Med. Best Med. Best Med. Best Med. Best Med. Best Med.
+Best-of-N $39.1$ $34.9$ $1.49$ $0.59$ $39.0$ $36.0$ $1.72$ $1.38$ $38.8$ $35.9$ $1.67$ $1.25$
+Zero–Order $39.5$ $36.4$ $1.54$ $0.98$ $39.4$ $37.4$ $1.75$ $1.60$ $38.9$ $36.1$ $1.67$ $1.29$
+Ours: Direct Noise Search
+CoSyNE $\mathbf{40.7}$ $\mathbf{38.3}$ $\mathbf{1.69}$ $\mathbf{1.47}$ $\mathbf{40.6}$ $\mathbf{39.0}$ $\mathbf{1.82}$ $\mathbf{1.75}$ $\mathbf{39.2}$ $\mathbf{36.8}$ $\mathbf{1.69}$ $\mathbf{1.45}$
+SNES $40.1$ $37.6$ $1.57$ $1.25$ $39.9$ $38.1$ $1.77$ $1.68$ $38.9$ $35.8$ $1.68$ $1.21$
+PGPE $39.0$ $36.1$ $1.37$ $0.88$ $38.4$ $36.4$ $1.67$ $1.50$ $38.7$ $35.8$ $1.65$ $1.23$
+[…]
+By virtue of our results, we show that optimizing noise with EAs is often sufficient to perform alignment (see Tab.˜1), though its effectiveness varies by model (Tab.˜8).
+This raises two questions: (1) Are some latent spaces easier to search than others?
+(2) How can one design, train, or modify diffusion models to make their latent spaces easier to search?
+Our reward improvements on PixArt-$\alpha$ were diminished compared to other diffusion models (Tab.˜8), suggesting some property of the latent space and/or diffusion model that affects our search efficacy.
+D.2 Effect of Longer Optimization Horizons on DNO
+Figures˜8(a), 8(b) and 8(c) depict the effect of longer optimization timelines on the best-sample, mean, and median rewards for the gradient-based method DNO, compared to gradient-free methods.
 Gradient-based require longer optimization timelines to reach the same reward as gradient-free methods.
-Figure 7: DNO with Long Optimization Horizon: Best-sample reward per step measured (ImageReward) as measured on Open Image Preferences.
-Horizontal lines are the maximum reward achieved by the gradient-free methods in 50 steps.
-Figure 8: DNO with Long Optimization Horizon: Mean reward per step measured (ImageReward) as measured on Open Image Preferences.
-Horizontal lines are the maximum (mean) reward achieved by the gradient-free methods in 50 steps.
-Figure 9: DNO with Long Optimization Horizon: Median reward per step measured (ImageReward) as measured on Open Image Preferences.
-Horizontal lines are the maximum (median) reward achieved by the gradient-free methods in 50 steps.
 D.3 Population Size Ablation
-Figures 10, 11 and 12 depict the effect of increasing population size on the CoSyne, SNES, and PGPE algorithms.
+Figures˜9, 10 and 11 depict the effect of increasing population size on the CoSyNE, SNES, and PGPE algorithms.
 All algorithms benefit from larger population sizes, however PGPE appears to benefit the most.
+D.4 Selection Pressure and Convergence
+Fig.˜12 shows the effect on increasing selection pressure (tournament size) on the median reward and standard deviation.
+We note that high selection pressure (larger tournament size) results in a rapid reduction in the reward standard deviation as shown in Fig.˜12(b), this is in line with prior work characterizing the effect of tournament selection [miller1995genetic].
+(a) Best-sample reward per step. Horizontal lines show the max reward in 50 steps.
+(b) Mean reward per step. Horizontal lines show the max mean in 50 steps.
+(c) Median reward per step. Horizontal lines show the max median in 50 steps.
+Figure 8: DNO with Long Optimization Horizon. ImageReward per step on Open Image Preferences.
 (a) Best-sample reward per step.
@@
 (c) Median reward per step.
-Figure 10: Genetic Algorithm: CoSyNE Reward statistics for the CosSyNE algorithm across optimization step.
+Figure 9: Genetic Algorithm: CoSyNE Reward statistics for the CoSyNE algorithm across optimization step.
 (a) Best-sample reward per step.
@@
 Figure 11: Evolutionary Strategy: PGPE Reward statistics for the PGPE algorithm across optimization step.
-D.4 Selection Pressure and Convergence
-Fig.
-13 shows the effect on increasing selection pressure (tournament size) on the median reward and standard deviation.
-We note that high selection pressure (larger tournament size) results in a rapid reduction in the reward standard deviation as shown in Fig.
-13(b), this is in line with prior work characterizing the effect of tournament selection [30].
 (a) Median reward per step.
@@
 Figure 12: Genetic Algorithm: CoSyNE Reward statistics per step, as measured on Open Image Preferences.
-Reward statistics are measured with increasing selection pressure, i.e., increasing tournament sizes of 2, 4, and 8.
+Reward statistics are measured with increasing selection pressure, i.e.
+increasing tournament sizes of 2, 4, and 8.
 We fix the population size to 16.
@@
 We then present sets of images that illustrate the low sample diversity of genetic algorithms across models and datasets.
-D.5.1 PGPE JPEG Reward-Hacking
-Fig.
-14 shows the proclivity of PGPE to "reward-hack" as noted in § 4.2.
+D.5.1 Reward Hacking in Sec.˜4.2
+Here we discuss reward-hacking behavior on our method, PGPE, and minor instances across other surveyed methods.
+Fig.˜17 shows the proclivity of PGPE to "reward-hack" as noted in Sec.˜4.2.
 In many instances, PGPE was able to reduce the JPEG file size, but the resulting images were far too dissimilar or nonsensical compared with the intended image for a prompt.
-This behavior is reward-hacking, because the algorithm completely neglects the prompt, and instead produces nonsensical images to maximize reward.
-Figure 14: PGPE Reward Hacking: Randomly selected images created using inference-time alignment with the PGPE algorithm (transformation search) to minimize the JPEG size (DrawBench).
-PGPE neglects the prompt, and naively attempts to minimize JPEG file size by producing monochromatic or simple images.
+This behavior is reward-hacking, because the algorithm completely ignores key prompt details, and instead produces nonsensical images to maximize reward.
 D.5.2 Stable Diffusion Qualitative Results
-Here, we present sets of images from DrawBench and Open Image Preferences to illustrate qualitative results — Fig.
-15, Fig.
-16, Fig.
-17, and Fig.
-18.
-Specifically, we show how genetic algorithms such as CoSyne feature low sample diversity, even as optimization steps scale.
-This can also happen to ES methods (Fig.
-18, however we noted it was less likely to occur with ES.
-Over longer optimization horizons, ES was able to maximize rewards better than GA while having higher diversity ( § 4.4).
+Here we illustrate the differences between ES and GA in longer optimization horizons.
+We show reward statistics in Fig.˜13, Fig.˜14, Fig.˜15, and Fig.˜16.
+GAs like CoSyNE feature low sample diversity, even as optimization steps scale.
+This can also happen to ES methods (Fig.˜16, however we noted it was less likely to occur with ES.
+Over longer optimization horizons, ES was able to maximize rewards better than GA while having higher diversity (Sec.˜4.3).
 Figure 13: Qualitative Sample Diversity: Randomly selected DrawBench prompts evaluated on StableDiffusion-1.5 with ImageReward.
@@
 Across optimization steps, the diversity between samples quickly diminishes.
-We identify this as a consistent occurrence with genetic algorithms such as CoSyne.
+Genetic algorithms such as CoSyNE are particularly vulnerable to this phenomenon.
 Figure 14: Qualitative Sample Diversity: Randomly selected Open Image Preference prompts evaluated on StableDiffusion-1.5 with ImageReward.
 CoSyNE was used to perform alignment.
-Note the low sample diversity.
+Note the low sample diversity across optimization steps.
+We identify this as a consistent occurrence with genetic algorithms such as CoSyne.
 Figure 15: Qualitative Sample Diversity: Randomly selected Open Image Preference prompts evaluated on StableDiffusion-3 with ImageReward.
@@
 D.6 Extended DrawBench Results
+Here we include additional, granular results on the DrawBench dataset.
+We provide this data in Tab.˜10, Tab.˜11, Tab.˜12, and Tab.˜13.
+Specifically, we include additional statistics (min, max, median, mean, standard) for each reward function.
+Each subtable for a particular reward function displays the population-best rewards and median rewards, which were measured across all prompts of the DrawBench dataset.
+D.7 Total Reward Function Evaluations
+The cost of evaluating the reward function is significantly higher than the cost of sampling [uehara2025inference]e.g.
+human feedback.
+Therefore, we report the number of total reward function evaluations from Tabs.˜1 and 2, which we illustrate in Tab.˜9.
+The evaluation costs are relatively low for the alignment objectives in this work.
+Even so, a method that can achieve high rewards with fewer samples is desirable in any context.
+We show our methods took substantially fewer reward evaluations than similar works FKD, SVDD, and DSearch-R, while achieving higher aesthetic scores than all other methods.
+For clarity, we describe how we derived the evaluation counts that appear in  Tab.˜9, which are based on the configurations from Tab.˜7.
+Best-of-N, Zero-Order.  These methods were parameterized by $P=16$, run over 15 optimization steps for 240 total evaluations. Each sample in the population was evaluated once per optimization step.
+CoSyNE.
+CoSyNE generates an intermediate population size that is $1.5\times$ its input and output population size $P$, as mentioned in Sec.˜4.5.
+Given $P=16$, an effective population of $P=24$ is realized; over 15 optimization steps this leads to 360 total evaluations.
+FKS.  FKS resampled latents a total of 41 times (denoising step 5 until step 45) over a population of 128 particles for 5258 total evaluations. Each particle is evaluated once at each denoising step.
+SVDD.  SVDD maintains 120 candidate latent noise vectors over 49 denoising steps for 5880 total evaluations. Each candidate is evaluated once at each denoising step.
+DSearch-R.
+DSearch-R maintains 5 candidate latent noise vectors, and follows a schedule to further expand its search, yielding $\sim$7880 evaluations under their default (exponential) search schedule.
+This resamples latents at $\sim$38 denoising steps.
+Figure 17: PGPE Reward Hacking: Randomly selected images created using inference-time alignment with the PGPE algorithm (transformation search) to minimize the JPEG size (DrawBench).
+PGPE neglects the prompt, and naively attempts to minimize JPEG file size by producing monochromatic or simple images.
+Table 9: Number of reward function evaluations per-prompt performed by each method from Tabs.˜1 and 2. If a reward evaluation was batched ($B>1$), we count that as $B$ evaluations.
+   Algorithm    Reward Evaluations
+   Best-of-N    240
+   Zero-Order    240
+   FKS    5258
+   SVDD    5880
+   DSearch-R    7880
+   CoSyNE    360
 Table 10: Extended DrawBench results with ImageReward.
@@
 Cosyne Rotation $0.299\pm 0.017$ 0.300 0.262 0.340
-SNES Noise $0.286\pm 0.016$ 0.288 0.247 0.324
+SNES Noise $6\pm 0.016 0.288 0.2$47 0.324
 SNES Rotation $0.290\pm 0.017$ 0.293 0.243 0.333
@@
 Random - $32.867\pm 3.502$ 32.812 20.562 41.750
-ZeroOrder - $34.545\pm 3.698$ 34.438 24.578 44.938
+ZeroOrder - $545\pm 3.698 34.43$8 24.578 44.938
 DNO - $26.237\pm 3.638$ 26.309 18.047 35.562
@@
 SNES Noise $35.623\pm 3.563$ 35.328 24.547 44.969
-SNES Rotation $34.699\pm 3.662$ 34.438 26.594 45.812
+SNES Rotation $699\pm 3.662 34.43$8 26.594 45.812
 PGPE Noise $34.397\pm 3.455$ 34.031 25.859 43.469
@@
 ZeroOrder - $53.35\pm 18.42$ 51.47 15.89 116.90
-DNO - $94.64\pm 22.78$ 91.78 47.11 174.45
 Cosyne Noise $39.52\pm 15.88$ 36.79 13.78 79.34
@@
 PGPE Rotation $20.18\pm 13.10$ 16.60 5.38 94.78
-Here, we include additional results on the DrawBench dataset.
-Specifically, we include additional statistics (min, max, median, mean, standard) for each reward function.
-Each subtable for a particular reward function displays the population-best rewards and median rewards, which were measured across all prompts of the DrawBench dataset.
-We provide this data in Tab.
-4, Tab.
-5, Tab.
-6, and Tab.
-7.